### v2

# arXiv:2505.19276v2[q-fin.RM] 03 Mar 2026

Abstract
We introduce a new paradigm for risk sharing that generalizes earlier models based on discrete agents and extends them to allow for sharing risk within a continuum of agents. Agents are represented by points of a measure space and have potentially heterogeneous risk preferences modeled by risk measures on a separable probability space. We derive the dual representation of the value function using a Strassen-type theorem for the weak-star topology and provide a characterization of the acceptance set using Aumann integration. These results are illustrated by explicit formulas when risk preferences are within the family of entropic and expected shortfall risk measures, and applications to Pareto efficiency in large markets.
1. Introduction
A significant literature studies the risk sharing problem: can one distribute a risk among finitely many agents, such that the total risk is minimized? Mathematically, for a given loss $\mathcal{X}$, this corresponds to the optimization problem
$\sum_{a\in A}\varrho_{a}(X_{a})\longrightarrow\mathrm{min!}$ (1)
subject to $\sum_{a\in A}X_{a}=\mathcal{X}$, where $(X_{a})_{a\in A}$ is a potential allocation of risk, modeled as a family of bounded random variables on a probability space $(\Omega,\mathscr{F},\mathbb{P})$. In (1), the finite set $A$ represents the space of agents, and is the index set for a collection $(\varrho_{a})_{a\in A}$ of risk measures, with $\varrho_{a}$ defining the preferences of agent $a$.
The risk sharing problem has a myriad of applications, including modern regulatory practice, where dispersing risk optimally within a market is often the ultimate goal. Optimal risk sharing is therefore of practical importance to policymakers. In particular, optimal risk sharing plays an important role in the theoretical underpinnings of capital requirements and capital adequacy tests (see [19]), and has practical consequences for the Solvency II directive regulating European insurance firms (see [28, 10]). Furthermore, firms subject to regulations may seek to avoid requirements by dividing their assets, leading to a minimization problem of the same form as (1) (see [27]). In a more indirect manner, solutions to the risk sharing problem characterize Pareto efficiency (see [14]).
However risk sharing is applied, the risk sharing framework implicit in (1) is not flawless; since $A$ is finite, each agent has a non-negligible impact on the model. Though appropriate for “too big to fail” banks, such impacts ignore the diffuse nature of smaller financial institutions, whether it be community banks, credit unions, or individual investors. These actors may have diverse considerations and preferences most accurately modeled—at least, as an approximation—by continuum models, where any one actor has a negligible impact on the model. Continuum approximations have precedence in the economics and game theory literature (see, for instance, the seminal paper of Aumann [4], or the theory of mean field games [6]).
While the above concerns justify including continuum agent models, they do not justify excluding all aspects of discrete agent spaces. In particular, the era of “too big to fail” has revealed the outsize influence of select financial institutions on the global market, even if sometimes counterbalanced by the combined impetus of smaller actors. Thus, it is necessary to consider both discrete and continuous agent spaces, potentially at the same time, reflecting the fact that some agents have essentially no impact on the market, while others may have disproportionate influence.
We therefore adopt an arbitrary finite complete measure space $(A,\mathscr{A},\mu)$ as an agent space, where finiteness refers to the assumption that $0<\mu(A)<\infty$. The measure $\mu$ may be purely atomic (corresponding to a discrete agent space), non-atomic (corresponding to a continuum of agents), or a mix between the two—allowing one to model a wide range of circumstances.
The cost of this universality is an increase in mathematical technicality. Under some circumstances, some allocations of risk must be excluded for failing to satisfy measurability. Furthermore, since spaces of random variables are often infinite dimensional, there are multiple ways to choose a notion of measurability for allocations even for a fixed $\sigma$-algebra on $A$. Nor is integration, the replacement of the finite sums in (1) and the associated constraint, easily assimilated into the theory; infinite dimensional spaces often support multiple versions of the classical Lebesgue integral. We adjudicate each of these issues, establishing a unified mathematical framework to answer the problems of risk sharing when agents form a general measure space.
Allowable allocations $(X_{a})_{a\in A}$ are assumed to be measurable, in a notion of measurability derived from the weak-star topology on $L^{\infty}(\mathbb{P})$, and integrable in the sense of Gelfand (see Definition 2). Risk preferences are represented by a collection of risk measures $(\varrho_{a})_{a\in A}$, which must also satisfy a measurability condition. More precisely, for every measurable allocation $(X_{a})_{a\in A}$, the mapping $a\longmapsto\varrho_{a}(X_{a})$ must be measurable. Once such assumptions are made, and an initial risk $\mathcal{X}\in L^{\infty}(\mathbb{P})$ is fixed, the minimization problem (1) can be stated as
$\int_{A}\varrho_{a}(X_{a})\mu(da)\longrightarrow\mathrm{min!}$ (2)
subject to the Gelfand integral of $(X_{a})_{a\in A}$ existing and equaling $\mathcal{X}$, potentially in addition to some other constraints.
If (2) is considered without constraints, we derive an explicit expression for the convex conjugate of the value function under general assumptions (see Theorem 1 and Appendix B). The formulas mimic the discrete case, replacing sums with integrals:
Theorem.
Suppose that the minimization problem (2) is considered without additional constraints and $\varrho(\mathcal{X})$ denotes the value function of the problem. If $(\varrho_{a})_{a\in A}$ are risk measures satisfying the Lebesgue property, $(\varrho_{a})_{a\in A}$ satisfies a measurability condition, $\int_{A}|{\varrho_{a}(0)}|\mu(da)<\infty$, and the value function is finite for all $\mathcal{X}$, then the value function is a risk measure, and the convex conjugate $\varrho^{\ast}$ satisfies
$\varrho^{\ast}(\mathbb{Q})=\int_{A}\varrho_{a}^{\ast}(\mathbb{Q})\mu(da)$
for any probability measure $\mathbb{Q}\ll\mathbb{P}$, where $\varrho^{\ast}_{a}$ denotes the convex conjugate of $\varrho_{a}$ for each $a\in A$. In particular,
$\varrho(\mathcal{X})=\sup_{\mathbb{Q}\ll\mathbb{P},\mathbb{Q}(\Omega)=1}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X})-\int_{A}\varrho_{a}^{\ast}(\mathbb{Q})\mu(da)\right).$
Proving the above formulas requires establishing a Strassen-type theorem for the weak-star topology, which we do in Appendix A.
As an illustration of the general theory developed for solving (2) in the unconstrained case, we give some concrete examples in §4, in particular for risk preferences within the family of entropic or expected shortfall risk measures. The formulas generalize previous results from the discrete case (including those of [9] and [24]). These are derived as a special case of formulas for dilations and inflations of a fixed risk measure. The former family is known and subsumes entropic risk measures, while the latter is a new definition, and includes expected shortfall as a special case. Although the risk sharing problem is always well-posed for dilated risk measures (see Theorem 2), we delineate sufficient conditions for the ill-posedness of the risk sharing problem for inflations of a fixed risk measure (see Theorem 4 and Appendix C).
In §5, we show the problem (2) is related to Pareto efficiency. More precisely, if the value function is finite, an allocation $(X_{a})_{a\in A}$ of $\mathcal{X}$ is Pareto efficient if, and only if, $(X_{a})_{a\in A}$ is a solution to the problem (2). As a consequence, the ill-posedness results from §4 imply economies where agents have preferences given by expected shortfall rarely have Pareto efficient allocations if there are a continuum of agents.
2. Notation and Preliminaries
Fix a complete measure space $(A,\mathscr{A},\mu)$, where $0<\mu(A)<\infty$. $A$ is the agent space, and elements $a\in A$ are agents. The spaces $L^{1}(\mu)$ and $L^{\infty}(\mu)$ carry their usual meaning.
Example 1.
A typical model for agents in the risk sharing literature is $A=\{1,\dots,N\}$, where $N\in\mathbb{N}$ is the number of agents. Equipping $A$ with the $\sigma$-algebra $2^{\{1,\dots,N\}}$ and the counting measure assimilates this model into our framework.∎
Example 2 (Aumann, [4]).
For $A=[0,1]$, let $\mathscr{A}$ be the Lebesgue $\sigma$-algebra on $A$, and let $\mu$ be the normalized Lebesgue measure on $\mathscr{A}$. Such a choice of the triple $(A,\mathscr{A},\mu)$ represents a continuum of agents, each with negligible individual impact on the model.∎
Example 3 (Shapley, [25]).
For $A=[0,1]$, let $\mathscr{A}$ be the Lebesgue $\sigma$-algebra on $A$. Denoting by $\lambda$ the normalized Lebesgue measure on $[0,1]$, define $\mu=\lambda+\delta_{0}+\delta_{1}$, where $\delta_{i}$ is the Dirac measure centered at $i$ ($i=0,1$). This corresponds to an agent space with two large agents, and infinitely many small agents, such that the combined force of the smaller agents is equal to half the combined force of the larger agents. ∎
To the author’s knowledge, neither of the agent spaces suggested by Example 2 or Example 3 have been considered by the risk measures literature before.
Each agent faces uncertainty, which is modeled by a separable probability space $(\Omega,\mathscr{F},\mathbb{P})$. The state of the world is completely described by a corresponding point $\omega\in\Omega$. $\mathscr{F}$ represents the amalgamation of information communicated about the state of the world by various observables.
The spaces $L^{1}(\mathbb{P})$ and $L^{\infty}(\mathbb{P})$ carry their usual meaning as spaces of contingent payoffs, although we adopt the convention that $\mathcal{X}\geq 0$ represents a loss of magnitude $\mathcal{X}$. $\mathscr{M}_{\mathbb{P}}$ will denote the set of absolutely continuous probability measures $\mathbb{Q}\ll\mathbb{P}$.
2.1. Allocations
It is necessary to consider payoffs parameterized by agents—viz., functions on $A$, taking values in $L^{\infty}(\mathbb{P})$. Such functions we call allocations. Applying an integration theory to such functions requires making measurability assumptions. To this end, let us introduce a notion of measurability.
Definition 1.
An allocation $(X_{a})_{a\in A}$ is said to be $\mathscr{A}$-measurable if, for each $\mathcal{Y}\in L^{1}(\mathbb{P})$, the function $a\longmapsto\mathbb{E}^{\mathbb{P}}(X_{a}\mathcal{Y})$ is $\mathscr{A}$-measurable. [^1] The above definition is equivalent to measurability with respect to the cylindrical $\sigma$-algebra on $L^{\infty}(\mathbb{P})$ generated by $L^{1}(\mathbb{P})\subseteq(L^{\infty}(\mathbb{P}))^{\ast}$, which is in turn equivalent to measurability with respect to the Baire $\sigma$-algebra of $\sigma(L^{\infty},L^{1})$ (see Theorem 2.3, [8]).
Equipped with the above notion, we may define an integration theory for allocations.
Definition 2.
An $\mathscr{A}$-measurable allocation $(X_{a})_{a\in A}$ is said to be Gelfand integrable if, for each $\mathcal{Y}\in L^{1}(\mathbb{P})$, the $\mathscr{A}$-measurable function $a\longmapsto\mathbb{E}^{\mathbb{P}}(X_{a}\mathcal{Y})$ is $\mu$-integrable.
If $(X_{a})_{a\in A}$ is Gelfand integrable, for each $B\in\mathscr{A}$, there exists a unique element $\mathcal{Z}_{B}\in L^{\infty}(\mathbb{P})$ such that
$\mathbb{E}^{\mathbb{P}}\left(\mathcal{Z}_{B}\mathcal{Y}\right)=\int_{B}\mathbb{E}^{\mathbb{P}}\left(X_{a}\mathcal{Y}\right)\mu(da),$
for each $\mathcal{Y}\in L^{1}(\mathbb{P})$ (see pg. 430, [1]). $\mathcal{Z}_{B}$ is called the Gelfand integral of $(X_{a})_{a\in A}$ over $B$, and is denoted $\int_{B}X_{a}\mu(da)$.
Subsequently, we consider the problem of distributing a risk $\mathcal{X}\in L^{\infty}(\mathbb{P})$ among the agent space $A$. The above integration theory allows us to formalize the allowable distributions.
Definition 3.
An $\mathscr{A}$-measurable allocation $(X_{a})_{a\in A}$ is said to be $\mathcal{X}$-feasible if $(X_{a})_{a\in A}$ is Gelfand integrable, and $\mathcal{X}=\int_{A}X_{a}\mu(da)$. The set of $\mathcal{X}$-feasible allocations is denoted $\mathbb{A}(\mathcal{X})$.
If $A=\{1,\dots,N\}$, all singletons are $\mathscr{A}$-measurable, and $1=\mu(\{1\})=\dots=\mu(\{N\})$ as in Example 1, $\mathcal{X}$-feasibility reduces to $\mathcal{X}=\sum_{a\in A}X_{a}$.
2.2. Risk Preferences
Each agent has risk preferences, which are modeled by a risk measure.
Definition 4.
A risk measure is a functional $\varrho:L^{\infty}(\mathbb{P})\longrightarrow\mathbb{R}$ satisfying properties (1) to (4) below. A risk measure $\varrho$ is said to have the Lebesgue property if it satisfies (5).
  1. 1.
    Monotonicity: for each $\mathcal{X},\mathcal{Y}\in L^{\infty}(\mathbb{P})$, if $\mathcal{X}\geq\mathcal{Y}$, then $\varrho(\mathcal{X})\geq\varrho(\mathcal{Y})$.
  2. 2.
    Cash additivity: for each $\mathcal{X}\in L^{\infty}(\mathbb{P})$, if $a\in\mathbb{R}$, then $\varrho(\mathcal{X}+a)=\varrho(\mathcal{X})+a$.
  3. 3.
    Convexity: for each $\mathcal{X},\mathcal{Y}\in L^{\infty}(\mathbb{P})$ and $\lambda\in[0,1]$, $\varrho(\lambda\mathcal{X}+(1-\lambda)\mathcal{Y})\leq\lambda\varrho(\mathcal{X})+(1-\lambda)\varrho(\mathcal{Y})$.
  4. 4.
    Fatou property: if $(\mathcal{X}^{n})_{n=1}^{\infty}\subseteq L^{\infty}(\mathbb{P})$ is an $L^{\infty}(\mathbb{P})$-bounded sequence converging in probability to $\mathcal{X}\in L^{\infty}(\mathbb{P})$, then
    $\varrho(\mathcal{X})\leq\liminf_{n\to\infty}\varrho(\mathcal{X}^{n}).$
  5. 5.
    Lebesgue property: if $(\mathcal{X}^{n})_{n=1}^{\infty}\subseteq L^{\infty}(\mathbb{P})$ is an $L^{\infty}(\mathbb{P})$-bounded sequence converging in probability to $\mathcal{X}\in L^{\infty}(\mathbb{P})$, then $\lim_{n\to\infty}\varrho(\mathcal{X}^{n})$ exists and equals $\varrho(\mathcal{X})$.
Risk measures are often defined to only satisfy (1) and (2) from Definition 4. Our definition is stronger, requiring (3) and (4) in addition to (1) and (2).
If $\varrho$ is a risk measure (not necessarily with the Lebesgue property), we have the dual representation
$\varrho(\mathcal{X})=\sup_{\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X})-\varrho^{\ast}(\mathbb{Q})\right),$ (3)
for each $\mathcal{X}\in L^{\infty}(\mathbb{P})$, where $\varrho^{\ast}(\mathbb{Q})=\sup_{\mathcal{X}\in L^{\infty}(\mathbb{P})}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X})-\varrho(X)\right)$ for each $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$. The function $\varrho^{\ast}$ is called the convex conjugate of $\varrho$, and is well-defined even if $\varrho$ is not a risk measure. If $\varrho^{\ast}(\mathscr{M}_{\mathbb{P}})\subseteq\{0,\infty\}$, we say that $\varrho$ is coherent.
For each agent $a\in A$, we therefore have a risk measure $\varrho_{a}$, codifying the risk preferences of agent $a$. Collecting all of the preferences yields a collection $(\varrho_{a})_{a\in A}$ of risk measures.
Consider now an $\mathcal{X}$-feasible allocation $(X_{a})_{a\in A}$. The goal of risk sharing is to minimize some measure of total risk $\mathrm{TR}$. Translating the formulas from the discrete case into the language of integration yields a formula of the form
$\mathrm{TR}=\int_{A}\varrho_{a}(X_{a})\mu(da).$
Unfortunately, the above integral need not be well-defined—it is unclear that the real-valued function $a\longmapsto\varrho_{a}(X_{a})$ is measurable or integrable. The integrability issue is settled by setting $\int_{A}\varrho_{a}(X_{a})\mu(da)=\infty$ whenever $a\longmapsto\varrho_{a}(X_{a})$ is measurable and $\int_{A}0\vee\varrho(X_{a})\mu(da)=\infty$. The measurability issue is resolved by restricting the possible collections of preferences $(\varrho_{a})_{a\in A}$ to those that satisfy the following definition.
Definition 5.
An indexed collection $(\varrho_{a})_{a\in A}$ of risk measures is said to be $\mathscr{A}$-measurable if, for each $\mathscr{A}$-measurable allocation $(X_{a})_{a\in A}$, the real-valued function $a\longmapsto\varrho_{a}(X_{a})$ is $\mathscr{A}$-measurable.
Example 4.
Since $(\Omega,\mathscr{F},\mathbb{P})$ is a separable probability space, the collection $(\varrho_{a})_{a\in A}$ defined by setting $\varrho_{a}=\varrho$ for all $a\in A$ for some fixed risk measure $\varrho$ is $\mathscr{A}$-measurable. [^2] Indeed, every convex lower $\sigma(L^{\infty},L^{1})$-semicontinuous function $\varphi$ can be represented as a supremum $\varphi(\mathcal{X})=\sup_{(\mathcal{Y},b)\in C}\left(\mathbb{E}^{\mathbb{P}}\left(\mathcal{Y}\mathcal{X}\right)+b\right)$ for all $\mathcal{X}\in L^{\infty}(\mathbb{P})$, where $C\subseteq L^{1}(\mathbb{P})\oplus\mathbb{R}$ is a set of $\sigma(L^{\infty},L^{1})$-continuous affine functions, which one can replace by a countable dense subset $C^{\prime}\subseteq C$ by virtue of separability, yielding measurability.
Example 5.
As a consequence of the conclusion of Example 4, the collection $(\varrho_{a})_{a\in A}$ defined by
$a\longmapsto\varrho_{a}=\sum_{i=1}^{n}\mathbf{1}_{B_{i}}(a)\varrho_{i}$
is $\mathscr{A}$-measurable for risk measures $\{\varrho_{1},\dots,\varrho_{n}\}$ and a disjoint $\mathscr{A}$-measurable partition $\{B_{1},\dots,B_{n}\}$ of AA.∎
Since many other preferences are simply a limiting case of Example 5, $\mathscr{A}$-measurability of $(\varrho_{a})_{a\in A}$ is not a stringent condition (see, in particular, Theorem 2 and Theorem 3 below).
2.3. The Risk Sharing Problem
Consider an element $\mathcal{X}\in L^{\infty}(\mathbb{P})$, to be allocated by a social planner among the agents in $A$, and fix an $\mathscr{A}$-measurable collection $(\varrho_{a})_{a\in A}$ of risk measures. The goal of the social planner is
$\int_{A}\varrho_{a}(X_{a})\mu(da)\longrightarrow\mathrm{min!}$
for $\mathcal{X}$-feasible allocations $(X_{a})_{a\in A}$ in a subset $C\subseteq\mathbb{A}(\mathcal{X})$. The subset $C$ can be strict, corresponding to a constrained version of the risk-sharing problem, or $C$ may equal all of $\mathbb{A}(\mathcal{X})$, corresponding to an unconstrained version of the risk-sharing problem. We consider only the unconstrained case in this article.
3. The Unconstrained Value Function
In this section, we study the value function of the risk sharing problem, which we introduce in §3.1. Our main result, Theorem 1, is contained in §3.2. Theorem 1 characterizes the convex conjugate of the value function, expressing it in terms of the the convex conjugates $(\varrho^{\ast}_{a})_{a\in A}$.
3.1. An Integral Infimal Convolution
Given risk measures $\varrho_{1}$ and $\varrho_{2}$, their infimal convolution $\varrho_{1}\operatorname{\Box}\varrho_{2}$ is defined by
$\left(\varrho_{1}\operatorname{\Box}\varrho_{2}\right)(\mathcal{X})=\inf_{\mathcal{Y}\in L^{\infty}(\mathbb{P})}\left(\varrho_{1}(\mathcal{Y})+\varrho_{2}(\mathcal{X}-\mathcal{Y})\right)$
for any $\mathcal{X}\in L^{\infty}(\mathbb{P})$. Naturally, the above definition can be extended to form the infimal convolution $\operatorname{\Box}_{i=1}^{N}\varrho_{i}$ of a finite set $\{\varrho_{1},\dots,\varrho_{N}\}$ of risk measures. Motivated by this definition, we define the integral infimal convolution, generalizing the classical concept, as follows.
Definition 6.
If $(\varrho_{a})_{a\in A}$ is an $\mathscr{A}$-measurable collection of risk measures, the integral infimal convolution $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ of $(\varrho_{a})_{a\in A}$ is defined as
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\inf_{(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})}\int_{A}\varrho_{a}(X_{a})\mu(da)$
for each $\mathcal{X}\in L^{\infty}(\mathbb{P})$.
[…]
It is not clear that $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ takes finite values. In fact, $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ can take the value -$\infty$. Since taking finite values is important for any application, we now note a sufficient condition for this to hold. Essentially, there must be at least partial agreement on priors. Such assumptions have appeared in the literature before to guarantee finiteness of the value function (see, for example, Condition (E) of [18]).
Proposition 1.
Suppose the $\mathscr{A}$-measurable collection $(\varrho_{a})_{a\in A}$ of risk measures consists of risk measures with the Lebesgue property, there exists $\mathbb{Q}\in\bigcap_{a\in A}\left\{\varrho^{\ast}_{a}<\infty\right\}$ such that $\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)<\infty$, and $\int_{A}|{\varrho_{a}(0)|}\mu(da)<\infty$. Then $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite: for each $\mathcal{X}\in L^{\infty}(\mathbb{P})$, we have that $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})<\infty$.
Note that, by Lemma 2 in Appendix A, Lemma 3 in Appendix B, and Lemma 4 in Appendix B (which assumes the Lebesgue property), the function $a\longmapsto\varrho^{\ast}_{a}(\mathbb{Q})$ is $\mathscr{A}$-measurable for each $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$. Thus, the integral $\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)$ is well-defined for each $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$ in the context of Proposition 1. At no other point in the proof of Proposition 1 do we use the Lebesgue property.
Proof.
We first show that $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})<\infty$ for all $\mathcal{X}\in L^{\infty}(\mathbb{P})$. Fixing an arbitrary $\mathcal{X}\in L^{\infty}(\mathbb{P})$, define $(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$ by $X_{a}=\mathcal{X}/\mu(A)$. Cash additivity implies
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})\leq\int_{A}\varrho_{a}(X_{a})\mu(da)\leq\int_{A}\left(|\varrho_{a}(0)|+\|{\mathcal{X}}\|_{L^{\infty}}/\mu(A)\right)\mu(da)<\infty,$
showing that $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})<\infty$. Since $\mathcal{X}\in L^{\infty}(\mathbb{P})$ was arbitrary, this proves the claim.
We now show that $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})>$-$\infty$ for all $\mathcal{X}\in L^{\infty}(\mathbb{P})$. Fix an arbitrary $\mathcal{X}\in L^{\infty}(\mathbb{P})$, and $\mathbb{Q}\in\bigcap_{a\in A}\left\{\varrho^{\ast}_{a}<\infty\right\}$ such that $\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)<\infty$. Then, for any $(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$,
$\int_{A}\varrho_{a}(X_{a})\mu(da)\geq\int_{A}\left(\mathbb{E}^{\mathbb{Q}}\left(X_{a}\right)-\varrho^{\ast}_{a}(\mathbb{Q})\right)\mu(da)=\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)-\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da).$
Thus, taking the infimum over $(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$,
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})\geq\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)-\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)>-\infty.$
Since $\mathcal{X}\in L^{\infty}(\mathbb{P})$ was arbitrary, this proves the claim. ∎
Even when $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite, it is not necessarily a risk measure, as it may fail to satisfy the Fatou property (an example is given for finite $A$ in [7]). Since this would prevent one from employing powerful duality techniques, the property of being a risk measure is a necessary assumption to make.
Although one can use duality theory for non-Fatou functionals, it requires employing finitely additive measures. For finite $A$, this causes no problems. However, with infinite $A$, measurability can become subtle, and the proper notion of measurability for allocations precludes the application of finitely additive measures to allocations, if measurability is to be preserved. Thus, compelled by necessity, we now consider sufficient conditions for the integral infimal convolution to possess the Fatou property.
Proposition 2.
Suppose the $\mathscr{A}$-measurable collection $(\varrho_{a})_{a\in A}$ of risk measures consists of risk measures with the Lebesgue property. Then, if $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite, $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is a risk measure.
Proof.
Monotonicity, convexity, and cash additivity are not difficult to prove, and we therefore focus on the Fatou property. Using a slight modification of the arguments in (Proposition 4.17, [11]), it suffices to prove continuity from above, in the sense that if $(\mathcal{X}^{n})_{n=1}^{\infty}\subseteq L^{\infty}(\mathbb{P})$ is decreasing and converges $\mathbb{P}$-a.s. to $\mathcal{X}\in L^{\infty}(\mathbb{P})$, then
$\inf_{n}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X}^{n})=\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X}).$
To this end, note that
$\inf_{n}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X}^{n})=\inf_{n}\inf_{(X_{a})_{a\in A}\in\mathbb{A}(0)}\int_{A}\varrho_{a}(\mathcal{X}^{n}/\mu(A)+X_{a})\mu(da)$
$\inf_{(X_{a})_{a\in A}\in\mathbb{A}(0)}\inf_{n}\int_{A}\varrho_{a}(\mathcal{X}^{n}/\mu(A)+X_{a})\mu(da)=\inf_{(X_{a})_{a\in A}\in\mathbb{A}(0)}\int_{A}\inf_{n}\varrho_{a}(\mathcal{X}^{n}/\mu(A)+X_{a})\mu(da)$
$\inf_{(X_{a})_{a\in A}\in\mathbb{A}(0)}\int_{A}\varrho_{a}(\mathcal{X}/\mu(A)+X_{a})\mu(da)=\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})$
by the monotone convergence theorem and the Lebesgue property of each $\varrho_{a}$, establishing the claim. ∎
3.2. Dual Representations
Given a finite set $\{\varrho_{1},\dots,\varrho_{N}\}$ of risk measures, it is known that their infimal convolution $\operatorname{\Box}_{i=1}^{N}\varrho_{i}$ satisfies
$\left(\operatorname{\Box}_{i=1}^{N}\varrho_{i}\right)^{\ast}=\sum_{i=1}^{N}\varrho^{\ast}_{i}.$ (4)
One can generalize this fact to the integral infimal convolution defined in the previous subsection. The formula remains essentially the same, although the finite sum in (4) is replaced by an integral.
Theorem 1.
Suppose the $\mathscr{A}$-measurable collection $(\varrho_{a})_{a\in A}$ of risk measures consists of risk measures with the Lebesgue property, $\int_{A}|{\varrho_{a}(0)}|\mu(da)<\infty$, and $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite. Then, for each $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$, the function $a\longmapsto\varrho^{\ast}_{a}(\mathbb{Q})$ is $\mathscr{A}$-measurable, and
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})=\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da).$ (5)
Furthermore,
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\sup_{\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X})-\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)\right).$ (6)
Proof.
See Appendix B for the proof of (5). Remark that, by Proposition 2, the integral infimal convolution is a risk measure under the assumptions of Theorem 1, and therefore (6) holds. ∎
In the process of proving Theorem 1, some interesting complementary results are derived in Appendix B. In particular, we characterize random variables with non-positive integral infimal convolution via Aumann integration (see Theorem 8), generalizing earlier results which used Minkowski summation (see, for example, the proof of Theorem 4.1, [20]).
4. Examples
We now apply the theory developed in §3 to specific families of risk measures, including those that fall in the class of entropic or expected shortfall risk measures.
In §4.1, dilations of a fixed risk measure $\varrho$ are considered, and explicit formulas for the value function and optimal allocation are given (see Theorem 2). Risk preferences modeled by entropic risk measures at various risk tolerance levels are covered by the results of this subsection.
In §4.2, we define a new family of risk measures obtained from a fixed coherent risk measure $\varrho$, which we call inflations of $\varrho$. An explicit formula is given for the value function when risk preferences are inflations of a fixed coherent risk measure (see Theorem 3), and sufficient conditions are given for the existence and non-existence of an optimal allocation (see Theorem 4 and Appendix C). Risk preferences modeled by expected shortfall at various quantile levels are covered by the results of this subsection.
4.1. Dilated Risk Measures
Given a risk measure $\varrho$ and $\gamma>0$, it is possible to construct a dilation $\varrho^{\gamma}$, which associates to $\varrho$ a family of risk measures. More precisely, the $\gamma$-dilation of $\varrho$ is the risk measure constructed by the following definition.
Definition 7.
Let $\varrho$ be a risk measure, and fix $\gamma>0$. The $\gamma$-dilation $\varrho_{\gamma}$ of $\varrho$ is defined by
$\varrho^{\gamma}(\mathcal{X})=\gamma\varrho\left(\frac{1}{\gamma}\mathcal{X}\right)$
for any $\mathcal{X}\in L^{\infty}(\mathbb{P})$.
In some circumstances, dilation may fail to produce any new non-trivial risk measures, as the following example illustrates.
Example 6.
Suppose $\varrho$ is a coherent risk measure, so that $\varrho^{\ast}(\mathscr{M}_{\mathbb{P}})\subseteq\{0,\infty\}$. Then, $\varrho^{\gamma}=\varrho$ for each $\gamma>0$. ∎
The triviality of Example 6 is not universal; in general, dilation can produce a non-trivial new family of risk measures, as demonstrated by the class of entropic risk measures.
Example 7.
For a risk tolerance parameter $\gamma>0$, the entropic risk measure $\mathrm{Ent}^{\gamma}$ is defined as
$\mathrm{Ent}^{\gamma}(\mathcal{X})=\gamma\log\left(\mathbb{E}^{\mathbb{P}}\left(e^{\frac{1}{\gamma}X}\right)\right)=\sup_{\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}}\left(\mathbb{E}^{\mathbb{Q}}(X)-\gamma D_{KL}(\mathbb{Q}\parallel\mathbb{P})\right)$
for any $\mathcal{X}\in L^{\infty}(\mathbb{P})$, where $D_{KL}(\mathbb{Q}\parallel\mathbb{P})=\mathbb{E}^{\mathbb{P}}\left(\frac{d\mathbb{Q}}{d\mathbb{P}}\log\left(\frac{d\mathbb{Q}}{d\mathbb{P}}\right)\right)$ is the Kullback-Leibler divergence. It is easy to see that $\mathrm{Ent}^{\gamma}$ is the $\gamma$-dilation of $\mathrm{Ent}^{1}$, and that $\mathrm{Ent}^{\gamma}\neq\mathrm{Ent}^{\gamma^{\prime}}$ for $\gamma\neq\gamma^{\prime}$ whenever $(\Omega,\mathscr{F},\mathbb{P})$ is sufficiently non-trivial. ∎
We now state the main result of this section, which explicitly derives the value function and optimal allocation when risk preferences are dilations of a fixed risk measure. Of particular note is the fact that an optimal allocation always exists, and an explicit formula is given. This explicit formula generalizes results from the discrete case (for instance, those of Righi and Moresco [24]).
Theorem 2.
Let $\varrho$ be a risk measure with the Lebesgue property, and let $(\gamma_{a})_{a\in A}\in(0,\infty)^{A}$ be an $\mathscr{A}$-measurable map with $\int_{A}\gamma_{a}\mu(da)<\infty$. Defining $\varrho_{a}=\varrho^{\gamma_{a}}$ for each $a\in A$ and $\Gamma=\int_{A}\gamma_{a}\mu(da)$, we have the following.
  1. 1.
    The indexed collection $(\varrho_{a})_{a\in A}$ of risk measures is $\mathscr{A}$-measurable.
  2. 2.
    The integral infimal convolution $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ satisfies
    $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)=\varrho^{\Gamma}.$
    In particular,
    $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}=\Gamma\varrho^{\ast}.$
  3. 3.
    For any $\mathcal{X}\in L^{\infty}(\mathbb{P})$, the allocation $(\gamma_{a}\mathcal{X}/\Gamma)_{a\in A}\in\mathbb{A}(\mathcal{X})$ is optimal, in the sense that
    $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\int_{A}\varrho_{a}(\gamma_{a}\mathcal{X}/\Gamma)\mu(da).$
Proof.
For (1), we may find a sequence $\left((\gamma^{n}_{a})_{a\in A}\right)_{n=1}^{\infty}$ of $\mathscr{A}$-measurable simple functions, such that $\lim_{n\to\infty}\gamma_{n}=\gamma$ pointwise. By replacing $\gamma_{n}$ with $\gamma_{n}\vee\frac{1}{n}$ if necessary, we may assume that $\gamma_{n}$ takes values in $(0,\infty)$. By the argument in Example 5, for each $n$, the family $(\varrho^{n}_{a})_{a\in A}$ of risk measures defined by $\varrho^{n}_{a}=\varrho^{\gamma^{n}_{a}}$ for each $a\in A$ is $\mathscr{A}$-measurable. Thus, for each $\mathscr{A}$-measurable allocation $(X_{a})_{a\in A}$, $a\longmapsto\varrho^{n}_{a}(X_{a})$ is $\mathscr{A}$-measurable. As $n\to\infty$, the Lebesgue property of $\varrho$ implies that $\lim_{n\to\infty}\varrho^{n}_{a}(X_{a})$ exists and equals $\varrho_{a}(X_{a})$. Since pointwise limits of $\mathscr{A}$-measurable functions are $\mathscr{A}$-measurable, this implies that $a\longmapsto\varrho_{a}(X_{a})$ is $\mathscr{A}$-measurable. Since $(X_{a})_{a\in A}$ was an arbitrary $\mathscr{A}$-measurable allocation, this shows that $(\varrho_{a})_{a\in A}$ is an $\mathscr{A}$-measurable collection of risk measures, proving (1).
To establish (2), we use Theorem 1. First, one must show that the preconditions for Theorem 1 hold. Thus, one must establish the following:
  1. i.
    $\int_{A}|{\varrho_{a}}(0)|\mu(da)<\infty$.
  2. ii.
    The collection $(\varrho_{a})_{a\in A}$ consists of risk measures with the Lebesgue property.
  3. iii.
    The integral infimal convolution $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite.
By the definition of dilation,
$\int_{A}|{\varrho_{a}(0)|}\mu(da)=\int_{A}|{\gamma_{a}\varrho(0)|}\mu(da)=\Gamma|{\varrho(0)}|<\infty,$
implying (i). Since $\varrho$ has the Lebesgue property, and all dilations of $\varrho$ therefore have the Lebesgue property, (ii) holds. To establish (iii), it suffices to verify the preconditions of Proposition 1; (i) and (ii) are both preconditions (both of which we have already verified), and the only remaining precondition is the existence of $\mathbb{Q}\in\bigcap_{a\in A}\{\varrho^{\ast}_{a}<\infty\}$ with $\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)<\infty$. There exists $\mathbb{Q}\in\{\varrho^{\ast}<\infty\}$; since $\varrho^{\ast}_{a}(\mathbb{Q})=\gamma_{a}\varrho^{\ast}(\mathbb{Q})$, it follows that $\mathbb{Q}\in\bigcap_{a\in A}\{\varrho^{\ast}_{a}<\infty\}$. It is easy to see that
$\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)=\int_{A}\gamma_{a}\varrho^{\ast}(\mathbb{Q})\mu(da)=\Gamma\varrho^{\ast}(\mathbb{Q})<\infty,$
establishing (iii).
We now apply Theorem 1. By Theorem 1, for all $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$,
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})=\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)=\int_{A}\gamma_{a}\varrho^{\ast}(\mathbb{Q})\mu(da)=\Gamma\varrho^{\ast}(\mathbb{Q})$
implying $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}=\Gamma\varrho^{\ast}$, and hence also that $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)=\varrho^{\Gamma}$. Thus, (2) holds.
For (3), it suffices to show (using the explicit representation of the integral infimal convolution previously derived) that
$\varrho^{\Gamma}(\mathcal{X})=\int_{A}\varrho_{a}(\gamma_{a}\mathcal{X}/\Gamma)\mu(da).$
By the definition of dilation,
$\int_{A}\varrho_{a}(\gamma_{a}\mathcal{X}/\Gamma)\mu(da)=\int_{A}\gamma_{a}\varrho(\mathcal{X}/\Gamma)\mu(da)=\Gamma\varrho(\mathcal{X}/\Gamma)=\varrho^{\Gamma}(\mathcal{X}),$
as desired. ∎
As an illustration of the above result, we now apply Theorem 2 to the case of entropic risk measures.
Example 8.
Suppose $(\gamma_{a})_{a\in A}\in(0,\infty)^{A}$ is $\mathscr{A}$-measurable, and $\int_{A}\gamma_{a}\mu(da)<\infty$. Then, by virtue of Theorem 2, the risk preferences $(\varrho_{a})_{a\in A}$ defined by $\varrho_{a}=\mathrm{Ent}^{\gamma_{a}}$ are such that
$\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)=\mathrm{Ent}^{\Gamma},$
where $\Gamma=\int_{A}\gamma_{a}\mu(da)$. Thus, the integral infimal convolution of entropic risk measures is an entropic risk measure with the risk tolerance parameter defined by the total risk tolerance of agents in $A$.
By Theorem 2, for a given risk $\mathcal{X}\in L^{\infty}(\mathbb{P})$ to allocate, an optimal allocation of risk is $(\gamma_{a}\mathcal{X}/\Gamma)_{a\in A}$. Under this allocation, each agent $a\in A$ receives the portion of $\mathcal{X}$ defined by considering their proportion $\gamma_{a}/\Gamma$ of the total risk tolerance $\Gamma$. ∎
4.2. Inflated Risk Measures
In this subsection, we introduce a new class of risk measures derived from a fixed coherent risk measure. Essentially, one enlarges the class of probability measures for which the convex conjugate returns a finite value.
Let $\varrho$ be a risk measure with the dual representation
$\varrho(\mathcal{X})=\sup_{\mathbb{Q}\in\{\varrho^{\ast}<\infty\}}\mathbb{E}^{\mathbb{Q}}(\mathcal{X}),$ (7)
where we assume $\mathbb{P}\in\{\varrho^{\ast}<\infty\}$. Within the class of risk measures taking the above form, the set $\{\varrho^{\ast}<\infty\}$ uniquely determines the dual representation of $\varrho$, and is denoted $\mathscr{Q}(\varrho)$. Define
$\widetilde{\mathscr{Q}}(\varrho)=\left\{\mathcal{Y}\in L^{1}(\mathbb{P}):\exists\mathbb{Q}\in\mathscr{Q}(\varrho)\textrm{ such that }0\leq\mathcal{Y}\leq\frac{d\mathbb{Q}}{d\mathbb{P}}\right\}.$
Definition 8.
Let $\varrho$ be a risk measure with dual representation (1), and fix a risk aversion parameter $\gamma\geq 1$. The $\gamma$-inflation $\widetilde{\varrho}_{\gamma}$ of $\varrho$ is defined by
$\widetilde{\varrho}_{\gamma}(\mathcal{X})=\sup_{\mathbb{Q}\in\gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}}\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)$
for any $\mathcal{X}\in L^{\infty}(\mathbb{P})$.
The idea of the above definition is that the set of probability measures used to calculate the dual representation is inflated by a factor of $\gamma$. To the author’s knowledge, the above definition is new, although some of the families generated by this definition are known, including expected shortfall.
Example 9.
For a quantile level $0<\alpha\leq 1$, define the risk measure $\mathrm{ES}^{\alpha}$ by
$\mathrm{ES}^{\alpha}(\mathcal{X})=\sup_{\mathbb{Q}\in\mathscr{M}_{\mathbb{P}},\frac{d\mathbb{Q}}{d\mathbb{P}}\leq\frac{1}{\alpha}}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$
for each $\mathcal{X}\in L^{\infty}(\mathbb{P})$. $\mathrm{ES}^{\alpha}$ is called the expected shortfall at quantile level $\alpha$. It is not difficult to see that $\mathrm{ES}^{\alpha}$ is the $\gamma$-inflation of $\mathrm{ES}^{1}=\mathbb{E}^{\mathbb{P}}$ for $\gamma=\frac{1}{\alpha}$. ∎
Remark 1.
If it is necessary to stress the underlying probability measure $\mathbb{P}$ from which expected shortfall is calculated, we will denote $\mathrm{ES}^{\alpha}=\mathrm{ES}^{\alpha}_{\mathbb{P}}$. ∎
In some sense, $\gamma\longmapsto\mathrm{ES}^{\frac{1}{\gamma}}$ is the canonical example of $\gamma$-inflation, as many properties of general inflated risk measures can be deduced from the corresponding properties of expected shortfall. One such example is continuity of the map $\gamma\longmapsto\widetilde{\varrho}_{\gamma}(\mathcal{X})$ for fixed $\mathcal{X}$, which is reducible to the case of expected shortfall, as we now demonstrate.
Proposition 3.
Fix $\mathcal{X}\in L^{\infty}(\mathbb{P})$. The map $\gamma\longmapsto\widetilde{\varrho}_{\gamma}(\mathcal{X})$ is left continuous on $(1,\infty)$.
Proof.
It suffices to show that, for each $\gamma^{\prime}\in(1,\infty)$ and $\varepsilon>0$, there exists $1\leq\gamma<\gamma^{\prime}$ and $\mathbb{Q}\in\gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$ with
$\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)\geq\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})-\varepsilon.$
There exists $\mathbb{Q}_{1}\in\gamma^{\prime}\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$ with
$\mathbb{E}^{\mathbb{Q}_{1}}\left(\mathcal{X}\right)\geq\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})-\frac{\varepsilon}{3}.$
Since $\mathbb{Q}_{1}\in\gamma^{\prime}\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$, there exists $\mathbb{Q}_{2}\in\mathscr{Q}(\varrho)$ with $\mathbb{Q}_{1}\ll\mathbb{Q}_{2}$ such that $\frac{d\mathbb{Q}_{1}}{d\mathbb{Q}_{2}}\leq\gamma^{\prime}$. Thus,
$\mathrm{ES}^{q(\gamma^{\prime})}_{\mathbb{Q}_{2}}(\mathcal{X})\geq\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})-\frac{\varepsilon}{3},$
where $q(x)=\frac{1}{x}$. Since expected shortfall is a continuous and decreasing function of quantile level (see [15] for an alternate integral definition of expected shortfall, from which continuity easily follows), and $x\longmapsto q(x)$ is continuous and decreasing, there exists $1\leq\gamma<\gamma^{\prime}$ such that
$\mathrm{ES}^{q(\gamma)}_{\mathbb{Q}_{2}}(\mathcal{X})\geq\mathrm{ES}^{q(\gamma^{\prime})}_{\mathbb{Q}_{2}}(\mathcal{X})-\frac{\varepsilon}{3}.$
There exists a probability measure $\mathbb{Q}\in\mathscr{M}_{\mathbb{Q}_{2}}\subseteq\mathscr{M}_{\mathbb{P}}$ with $\frac{d\mathbb{Q}}{d\mathbb{Q}_{2}}\leq\gamma$ and
$\mathbb{E}^{\mathbb{Q}}(\mathcal{X})\geq\mathrm{ES}^{q(\gamma)}_{\mathbb{Q}_{2}}(\mathcal{X})-\frac{\varepsilon}{3}.$
Combining everything, we obtain that
$\mathbb{E}^{\mathbb{Q}}(\mathcal{X})\geq\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})-\varepsilon,$
which proves the claim, as $\frac{d\mathbb{Q}}{d\mathbb{P}}\leq\gamma\frac{d\mathbb{Q}_{2}}{d\mathbb{P}}$, implying $\mathbb{Q}\in\gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$. ∎
For a family $(\gamma_{a})_{a\in A}\in[1,\infty)^{A}$ and a risk measure $\varrho$ it is possible to construct a collection $(\varrho_{a})_{a\in A}$ of risk preferences via $\varrho_{a}=\widetilde{\varrho}_{\gamma_{a}}$. Given such a family, we now consider the value function of the risk sharing problem, characterizing the dual representation of the value function, and ensuring $(\varrho_{a})_{a\in A}$ satisfies the requisite measurability condition under broad circumstances.
Theorem 3.
Let $\varrho$ be a risk measure with the Lebesgue property and the representation (7). Let $(\gamma_{a})_{a\in A}\in[1,\infty)^{A}$ be an $\mathscr{A}$-measurable map, with $\mu$-essential infimum $\Gamma$. Defining $\varrho_{a}=\widetilde{\varrho}_{\gamma_{a}}$ for each $a\in A$, we have the following.
  1. 1.
    The indexed collection $(\varrho_{a})_{a\in A}$ of risk measures is $\mathscr{A}$-measurable.
  2. 2.
    The integral infimal convolution $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ satisfies
    $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)=\widetilde{\varrho}_{\Gamma}.$
    In particular,
    $\mathscr{Q}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)=\Gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}.$
Proof.
For (1), we may find an increasing sequence $\left((\gamma^{n}_{a})_{a\in A}\right)_{n=1}^{\infty}$ of $\mathscr{A}$-measurable simple functions, such that $\gamma_{n}\uparrow\gamma$ pointwise. Furthermore, we may assume that $\gamma_{n}$ takes values in $1,\infty)$ (indeed, one can replace $\gamma_{n}$ with the $\mathscr{A}$-measurable simple function $\gamma_{n}\vee 1$). By the argument in Example 5, for each $n$, the family $(\varrho^{n}_{a})_{a\in A}$ of risk measures defined by $\varrho^{n}_{a}=\widetilde{\varrho}_{\gamma^{n}_{a}}$ for each $a\in A$ is $\mathscr{A}$-measurable. Thus, for each $\mathscr{A}$-measurable allocation $(X_{a})_{a\in A}$, $a\longmapsto\varrho^{n}_{a}(X_{a})$ is $\mathscr{A}$-measurable. As $n\to\infty$, Proposition 3 implies that $\lim_{n\to\infty}\varrho^{n}_{a}(X_{a})$ exists and equals $\varrho_{a}(X_{a})$. Since pointwise limits of $\mathscr{A}$-measurable functions are $\mathscr{A}$-measurable, this implies that $a\longmapsto\varrho_{a}(X_{a})$ is $\mathscr{A}$-measurable. Since $(X_{a})_{a\in A}$ was an arbitrary $\mathscr{A}$-measurable allocation, this shows that $(\varrho_{a})_{a\in A}$ is an $\mathscr{A}$-measurable collection of risk measures, proving (1).
To prove (2), we apply Theorem 1. First, one must show that the preconditions for Theorem 1 hold. Thus, one must establish the following:
  1. i.
    $\int_{A}|{\varrho_{a}}(0)|\mu(da)<\infty$.
  2. ii.
    The collection $(\varrho_{a})_{a\in A}$ consists of risk measures with the Lebesgue property.
  3. iii.
    The integral infimal convolution $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite.
Clearly, since $\varrho_{a}(0)=0$ for all $a$, (i) holds. For (ii), note that the Jouini-Schachermayer-Touzi theorem (see Theorem 2.4, [21]) implies that, since $\varrho$ has the Lebesgue property, $\mathscr{Q}(\varrho)$ must be uniformly integrable (viewed as a subset of $L^{1}(\mathbb{P})$ via the Radon-Nikodým derivative). Thus, for each $\gamma^{\prime}\geq 1$, $\gamma^{\prime}\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$ is uniformly integrable. Since any risk measure representable as a supremum of expectations over a uniformly integrable set of probability measures has the Lebesgue property, it follows that every inflation of $\varrho$ has the Lebesgue property. In particular, $(\varrho_{a})_{a\in A}$ consists of risk measures with the Lebesgue property, and (ii) therefore holds. To establish (iii), it suffices to verify the preconditions of Proposition 1; (i) and (ii) are both preconditions (both of which we have already verified), and the only remaining precondition is the existence of $\mathbb{Q}\in\bigcap_{a\in A}\{\varrho^{\ast}_{a}<\infty\}$ with $\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da)<\infty$. For this last precondition, fix some $\mathbb{Q}\in\{\varrho^{\ast}<\infty\}\neq\emptyset$. For each $a\in A$, $\varrho^{\ast}_{a}(\mathbb{Q})=0$, proving the claim.
We now apply Theorem 1. It suffices to show that $\mathbb{Q}\in\Gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$ if, and only if, $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})<\infty$. If $\mathbb{Q}\in\Gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$, then $\varrho^{\ast}_{a}(\mathbb{Q})=0$ for $\mu$-a.e. $a\in A$, implying (via Theorem 1) that $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})<\infty$. Conversely, suppose $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})<\infty$. It is easy to see that $\mathbb{Q}\in\gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}$ for each $\gamma>\Gamma$. Thus,
$\mathbb{Q}\in\bigcap_{\gamma>\Gamma}\gamma\widetilde{\mathscr{Q}}(\varrho)\cap\mathscr{M}_{\mathbb{P}}=\mathscr{M}_{\mathbb{P}}\cap\bigcap_{\gamma>\Gamma}\gamma\widetilde{\mathscr{Q}}(\varrho),$
implying it suffices to show that $\Gamma\widetilde{\mathscr{Q}}(\varrho)=\bigcap_{\gamma>\Gamma}\gamma\widetilde{\mathscr{Q}}(\varrho)$. Fix any $\mathcal{X}\in\bigcap_{\gamma>\Gamma}\gamma\widetilde{\mathscr{Q}}(\varrho)$; it suffices to show that $\mathcal{X}\in\Gamma\widetilde{\mathscr{Q}}(\varrho)$. Take a strictly decreasing $(\Gamma_{n})_{n=1}^{\infty}\downarrow\Gamma$; for each $n$, we may find $\mathbb{Q}_{n}\in\mathscr{Q}(\varrho)$ with
$0\leq\mathcal{X}\leq\Gamma_{n}\frac{d\mathbb{Q}_{n}}{d\mathbb{P}}.$
As established before, $\mathscr{Q}(\varrho)$ is uniformly integrable. Thus, by Mazur’s lemma (see Theorem 3.19, [5]) and the Dunford-Pettis theorem, [^3] The Dunford-Pettis theorem asserts that a subset of $L^{1}(\mathbb{P})$ is relatively $\sigma(L^{1},L^{\infty})$-compact if, and only if, it is uniformly integrable. there exists $\widetilde{\mathbb{Q}}_{n}\in\mathrm{co}\{\mathbb{Q}_{m}:m\geq n\}$ such that $(\widetilde{\mathbb{Q}}_{n})_{n=1}^{\infty}$ converges to some $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$ in $L^{1}(\mathbb{P})$ (equivalently, in total variation norm); since $\mathscr{Q}(\varrho)$ is closed in $L^{1}(\mathbb{P})$, $\mathbb{Q}\in\mathscr{Q}(\varrho)$. It is easy to see that
$0\leq\mathcal{X}\leq\Gamma_{n}\frac{d\widetilde{\mathbb{Q}}_{n}}{d\mathbb{P}}.$
Thus, using Borel-Cantelli to pass to a $\mathbb{P}$-a.s. convergent subsequence if necessary, we have that
$0\leq\mathcal{X}\leq\Gamma\frac{d\mathbb{Q}}{d\mathbb{P}},$
showing that $\mathcal{X}\in\Gamma\widetilde{\mathscr{Q}}(\varrho)$, as desired. ∎
We consider now whether the infimum inherent in the value function is attained, in the setting of Theorem 3. There are two circumstances to consider, depending on the nature of the essential infimum $\Gamma$ of $(\gamma_{a})_{a\in A}$:
  1. 1.
    $\mu(\{a:\gamma_{a}=\Gamma\})>0$, in which case an optimal allocation is found by giving all the risk to the agents $a$ such that $\gamma_{a}=\Gamma$.
  2. 2.
    $\mu(\{a:\gamma_{a}=\Gamma\})=0$, in which case the existence of optimal allocations becomes subtle. Intuitively, the infimum should not be attained, since one should be able to shift risk from agents $a$ with $\gamma_{a}>\Gamma+\varepsilon$ (where $0<\varepsilon\ll 1$) to agents $b$ with $\gamma_{b}\leq\Gamma+\varepsilon$, constituting an improvement on an apparently optimal allocation (this intuition is formalized in Appendix C). However, the infimum is always attained if the risk $\mathcal{X}$ to be allocated is a constant random variable, or more generally if $\gamma\longmapsto\widetilde{\varrho}_{\gamma}(\mathcal{X})$ is constant on $\Gamma,\infty)$. Thus, to conclude an optimal allocation does not exist, one must introduce a condition on $\mathcal{X}$ ensuring it is not unaffected by a change in the risk aversion parameter $\gamma$.
Our main result in this direction is Theorem 4 below. Compared to the finite agent case for expected shortfall (see, for example, [9]), our result simultaneously exhibits new phenomena and generalizes known results: when $\mu(\{a:\gamma_{a}=\Gamma\})>0$, the finite agent formulas for an optimal allocation remain true, while if $\mu(\{a:\gamma_{a}=\Gamma\})=0$, an optimal allocation may fail to exist, something which is not true in the finite agent case.
Theorem 4.
Let $\varrho$ be a risk measure with the Lebesgue property and the representation (7). Let $(\gamma_{a})_{a\in A}\in[1,\infty)^{A}$ be an $\mathscr{A}$-measurable map, with $\mu$-essential infimum $\Gamma$. Defining $\varrho_{a}=\widetilde{\varrho}_{\gamma_{a}}$ for each $a\in A$, we have the following.
  1. 1.
    Suppose $\mu(\{a:\gamma_{a}=\Gamma\})>0$. Then, for any $\mathcal{X}\in L^{\infty}(\mathbb{P})$, the allocation
    $(\mathbf{1}_{\{b:\gamma_{b}=\Gamma\}}(a)\mathcal{X}/\mu(\{b:\gamma_{b}=\Gamma\}))_{a\in A}\in\mathbb{A}(\mathcal{X})$
    is optimal, in the sense that
    $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\int_{A}\varrho_{a}(\mathbf{1}_{\{b:\gamma_{b}=\Gamma\}}(a)\mathcal{X}/\mu(\{b:\gamma_{b}=\Gamma\}))\mu(da).$
  2. 2.
    Suppose the following conditions are true for $\mathcal{X}\in L^{\infty}(\mathbb{P})$.
    1. (a)
      $\mu(\{a:\gamma_{a}=\Gamma\})=0$.
    2. (b)
      There exists $\Gamma^{\prime}>\Gamma$ so that $\widetilde{\varrho}_{\Gamma}(\mathcal{X})\neq\widetilde{\varrho}_{\Gamma^{\prime}}(\mathcal{X})$.
    Then there does not exist an allocation $(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$ such that
    $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\int_{A}\varrho_{a}(X_{a})\mu(da).$
Proof.
We prove the first assertion here; the proof of the second assertion is contained in Appendix C.
For the first assertion, note that Theorem 3 implies equivalence to the claim that
$\widetilde{\varrho}_{\Gamma}(\mathcal{X})=\int_{A}\varrho_{a}(\mathbf{1}_{\{b:\gamma_{b}=\Gamma\}}(a)\mathcal{X}/\mu(\{b:\gamma_{b}=\Gamma\}))\mu(da).$
Clearly,
$\int_{A}\varrho_{a}(\mathbf{1}_{\{b:\gamma_{b}=\Gamma\}}(a)\mathcal{X}/\mu(\{b:\gamma_{b}=\Gamma\}))\mu(da)=\int_{\{b:\gamma_{b}=\Gamma\}}\frac{1}{\mu(\{b:\gamma_{b}=\Gamma\})}\widetilde{\varrho}_{\Gamma}\left(\mathcal{X}\right)\mu(da)$
$\widetilde{\varrho}_{\Gamma}(\mathcal{X}),$
as desired, proving the first assertion. ∎
5. Applications to Pareto Efficiency
As a consequence of Theorem 4, it is possible to show certain “simple” markets, constructed only using expected shortfall, do not admit Pareto efficient allocations if there are a continuum of agents.
We recall some terminology from economics.
Definition 9.
Let $(\varrho_{a})_{a\in A}$ be an $\mathscr{A}$-measurable collection of preferences. An $\mathcal{X}$-feasible allocation $(X_{a})_{a\in A}$ is said to be Pareto efficient if there does not exist another $\mathcal{X}$-feasible allocation $(Y_{a})_{a\in A}$ with $\varrho_{a}(Y_{a})\leq\varrho_{a}(X_{a})$ $\mu$-a.e. and $\mu\left(\left\{\varrho_{a}(Y_{a})<\varrho_{a}(X_{a})\right\}\right)>0$. If such a $(Y_{a})_{a\in A}$ exists for an $\mathcal{X}$-feasible $(X_{a})_{a\in A}$, $(Y_{a})_{a\in A}$ is said to be a Pareto improvement on $(X_{a})_{a\in A}$.
It is known Pareto efficiency in the discrete case is related to the optimization of a social welfare function via the Negishi weight method, [^4] In our setting, where we work with risk measures and not utility functions, it may be more appropriate to call the social welfare function a social loss/risk function. and that, under translation invariance of the preferences, this reduces to the classical risk sharing problem (see, for example, Proposition 3.3, [14]). Similarly, we have the following result, essentially due to Ghossoub and Nendel [13].
Proposition 4.
Suppose $(\varrho_{a})_{a\in A}$ is an $\mathscr{A}$-measurable collection of risk measures. If $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})<\infty$, then an $\mathcal{X}$-feasible allocation $(X_{a})_{a\in A}$ is Pareto efficient if, and only if,
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\int_{A}\varrho_{a}(X_{a})\mu(da).$
Proof.
If $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})=\int_{A}\varrho_{a}(X_{a})\mu(da)$, and there exists a Pareto improvement $(Y_{a})_{a\in A}$ on $(X_{a})_{a\in A}$, then
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})\leq\int_{A}\varrho_{a}(Y_{a})\mu(da)<\int_{A}\varrho_{a}(X_{a})\mu(da)=\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X}),$
a contradiction.
Conversely, suppose $(X_{a})_{a\in A}$ is Pareto efficient, but
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})<\int_{A}\varrho_{a}(X_{a})\mu(da).$
Hence, there exists $\mathcal{X}$-feasible $(Y_{a})_{a\in A}$ with
$\int_{A}\varrho_{a}(Y_{a})\mu(da)<\int_{A}\varrho_{a}(X_{a})\mu(da).$
Define $R=\frac{1}{\mu(A)}\int_{A}\left(\varrho_{a}(X_{a})-\varrho_{a}(Y_{a})\right)\mu(da)<0$. Consider $(Z_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$ defined as
$Z_{a}=Y_{a}+\varrho_{a}(X_{a})-\varrho_{a}(Y_{a})-R.$
For each $a\in A$, $\varrho_{a}(Y_{a}+\varrho_{a}(X_{a})-\varrho_{a}(Y_{a}))=\varrho_{a}(X_{a})$. Hence, as $R>0$, for each $a\in A$, $\varrho_{a}(Z_{a})=\varrho_{a}(X_{a})-R<\varrho_{a}(X_{a})$, showing $(Z_{a})_{a\in A}$ constitutes a Pareto improvement on $(X_{a})_{a\in A}$, contradicting the Pareto efficiency of $\mathcal{X}$. ∎
Remark 2.
The second part of the above argument is based on Kaldor-Hicks efficiency (for a definition and a discussion of Kaldor-Hicks efficiency, see §1.2, Posner [22]). More precisely, it shows that if $(Y_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$ has a lower total risk than $(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$, it is possible for the beneficiaries of the lower risk total from $(Y_{a})_{a\in A}$ to compensate with cash those that are left worse off from $(Y_{a})_{a\in A}$. ∎
Conjoining Proposition 4 with the ill-posedness results of §4, we obtain a non-existence result for Pareto optima.
Theorem 5.
Suppose $(\beta_{a})_{a\in A}\in(0,1]^{A}$ is an $\mathscr{A}$-measurable map with $\mu$-essential supremum $\beta>0$, and $\mathcal{X}\in L^{\infty}(\mathbb{P})$ is an aggregate endowment. Define the individual preferences $(\varrho_{a})_{a\in A}$ by $\varrho_{a}=\mathrm{ES}^{\beta_{a}}_{\mathbb{P}}$. If $\mu(\{a:\beta_{a}=\beta\})=0$ and there exists $0<\alpha<\beta$ so $\mathrm{ES}^{\beta}_{\mathbb{P}}(\mathcal{X})\neq\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})$, then there does not exist any Pareto efficient $\mathcal{X}$-feasible allocations.
Proof.
Theorem 4 implies there cannot exist $(X_{a})_{a\in A}\in\mathbb{A}(\mathcal{X})$ with
$\int_{A}\varrho_{a}(X_{a})\mu(da)=\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X}).$ (8)
Since any Pareto efficient allocation of $\mathcal{X}$ must satisfy (8) (see Proposition 4), it follows that Pareto efficient allocations of $\mathcal{X}$ cannot exist. ∎
Appendix A The Weak-Star Strassen Theorem
Theorem 2.2 of Hiai and Umegaki [16] establishes an integral exchange formula for correspondences valued in a separable Banach space, allowing one to swap an infimum (equivalently, a supremum) and an integral, allowing one to characterize the support function of certain set integrals. The Hiai-Umegaki result is an example of a Strassen-type theorem. Strassen-type theorems have been extended to cover correspondences valued in separable Banach spaces equipped with the weak topology (see [2]), correspondences valued in the compact subsets of a locally convex topological vector space (see [26]), and have ramifications even in the finite dimensional case (see [3]). However, the literature on this topic is both highly technical and likely not directly applicable to our circumstances (e.g., requiring the correspondence to take weak-star compact values). Thus, in this section, we derive a Strassen-type theorem for correspondences valued in the dual of a separable Banach space and taking weak-star closed values, with measurability and integration understood in a weak-star sense.
A.1. Notation
Let $(A,\mathscr{A},\mu)$ denote a finite complete measure space. The trace $\sigma$-algebra $\{C\cap B:C\in\mathscr{A}\}$ of $B\in\mathscr{A}$ is denoted $\mathscr{A}_{B}$. Let $E$ be a separable Banach space; denote by $B_{E}$ the unit ball of $E$, and let $B^{\ast}=B^{\circ}_{E}$ denote the closed unit ball of $E^{\ast}$, the dual of $E$.
A function $f:A\longrightarrow E^{\ast}$ is said to be $\mathscr{A}$-measurable if $a\longmapsto\langle{x,f(a)}\rangle$ is $\mathscr{A}$-measurable for each $x\in E$. An $\mathscr{A}$-measurable function $f:A\longrightarrow E^{\ast}$ is said to be Gelfand integrable if $\langle{x,f}\rangle\in L^{1}(\mu)$ for each $x\in E$. If $f$ is Gelfand integrable, and $B\in\mathscr{A}$, there exists a unique element $g_{B}\in E^{\ast}$ such that $\langle{x,g_{B}}\rangle=\int_{B}\langle{x,f(a)}\rangle\mu(da)$. The element $g_{B}$ is denoted $\int_{B}f(a)\mu(da)$, and is called the Gelfand integral of $f$ over $B$. These notions all parallel those introduced for $E^{\ast}=L^{\infty}(\mathbb{P})$ in §2.1.
Consider a correspondence $F:A\longrightarrow 2^{E^{\ast}}$. Given a subset $U\subseteq E^{\ast}$, define
$F^{-1}(U)=\{a\in A:F(a)\cap U\neq\emptyset\}.$
An integrable selector of $F$ is a Gelfand integrable function $f:A\longrightarrow E^{\ast}$ such that $f(a)\in F(a)$ for $\mu$-a.e. $a\in A$. The set of integrable selectors of $F$ is denoted $S^{1}(F)$.
Definition 10.
The Aumann integral of $F$, denoted $\int_{A}F(a)\mu(da)$, is the subset of $E^{\ast}$ defined by
$\int_{A}F(a)\mu(da)=\left\{\int_{A}f(a)\mu(da):f\in S^{1}(F)\right\}.$
In a similar fashion to the above concept of integration, one can introduce notions of measurability for correspondences.
Definition 11.
$F$ is said to be $\mathscr{A}$-measurable if $F^{-1}(U)\in\mathscr{A}$ for every $\sigma(E^{\ast},E)$-closed $U\subseteq E^{\ast}$.
Definition 12.
$F$ is said to be Effros $\mathscr{A}$-measurable if $F^{-1}(U)\in\mathscr{A}$ for every $\sigma(E^{\ast},E)$-open $U\subseteq E^{\ast}$.
Definition 11 and Definition 12 are coherent for correspondences valued in any topological space, not just $(E^{\ast},\sigma(E^{\ast},E))$. In the sequel, we generally employ $\mathscr{A}$-measurability rather than Effros $\mathscr{A}$-measurability, since the former has better stability properties. However, since many results are stated in terms of Effros $\mathscr{A}$-measurability, we cannot expunge Definition 12 from our analysis.
A.2. Preliminary Results
In this subsection, we use $F$ to denote an $\mathscr{A}$-measurable correspondence $A\longrightarrow 2^{E^{\ast}}$ with non-empty $\sigma(E^{\ast},E)$-closed values.
For $\lambda\geq 0$, define a correspondence $F_{\lambda}:A\longrightarrow 2^{\lambda B^{\ast}}$ by $F_{\lambda}=F\cap\lambda B^{\ast}$. $F_{\lambda}$ is $\mathscr{A}$-measurable. Define $\widetilde{F}_{\lambda}$ as the restriction of $F_{\lambda}$ to $R_{\lambda}=F^{-1}(\lambda B^{\ast})\in\mathscr{A}$; it is easy to see that $\widetilde{F}_{\lambda}$ is $\mathscr{A}_{R_{\lambda}}$-measurable.
Lemma 1.
There exists a collection $\{f_{n}:n\in\mathbb{N}\}\subseteq(E^{\ast})^{A}$ of $\mathscr{A}$-measurable functions such that
$F(a)=\overline{\bigcup_{n\in\mathbb{N}}\{f_{n}(a)\}}^{\sigma(E^{\ast},E)}$
for each $a\in A$.
Proof.
Take $\lambda$ large enough so $R_{\lambda}\neq\emptyset$. By a result of Himmelberg [17], since $\lambda B^{\ast}$ is a Polish space and $\widetilde{F}_{\lambda}$ is $\mathscr{A}_{R_{\lambda}}$-measurable, there exists a collection $\{g_{n}^{\lambda}:n\in\mathbb{N}\}\subseteq(E^{\ast})^{R_{\lambda}}$ of $\mathscr{A}_{R_{\lambda}}$-measurable functions such that
$\widetilde{F}_{\lambda}(a)=\overline{\bigcup_{n\in\mathbb{N}}\left\{g_{n}^{\lambda}(a)\right\}}^{\sigma(E^{\ast},E)}$
for each $a\in R_{\lambda}$. If there exists an $\mathscr{A}$-measurable $h:A\longrightarrow E^{\ast}$ such that $h(a)\in F(a)$ for each $a\in A$, the claim is proved. Indeed, take $(\lambda_{n})_{n=1}^{\infty}\uparrow\infty$, and consider the countable collection $\left\{k^{n}_{m}:(n,m)\in\mathbb{N}\times\mathbb{N}\right\}$ of $\mathscr{A}$-measurable functions defined by
$k^{n}_{m}|_{R_{\lambda_{n}}}=g^{\lambda_{n}}_{m},$
$k^{n}_{m}|_{A\setminus R_{\lambda_{n}}}=h.$
Then $F(a)=\overline{\bigcup_{(n,m)\in\mathbb{N}\times\mathbb{N}}\{k^{n}_{m}(a)\}}^{\sigma(E^{\ast},E)}$, proving the claim, since one can consider a bijection $\mathbb{N}\longrightarrow\mathbb{N}\times\mathbb{N}$.
We now prove the existence of an $\mathscr{A}$-measurable $h:A\longrightarrow E^{\ast}$ such that $h(a)\in F(a)$ for each $a\in A$. There exists a disjoint partition $\{D_{n}:n\in G\subseteq\mathbb{N}\}\subseteq\mathscr{A}\setminus\{\emptyset\}$ of $A$ such that for each $n\in G$, $D_{n}\subseteq F^{-1}(\lambda^{\prime}B^{\ast})$ for large enough $\lambda^{\prime}$ (which may depend on nn). [^5] For example, take $\widetilde{\lambda}$ large enough so that $R_{\widetilde{\lambda}}\neq\emptyset$. Define $D_{1}=F^{-1}\left(\widetilde{\lambda}B^{\ast}\right)$, and let $D_{n+1}=F^{-1}\left((n+1)\widetilde{\lambda}B^{\ast}\right)\setminus F^{-1}\left(n\widetilde{\lambda}B^{\ast}\right)$. Taking $G=\{n:D_{n}\neq\emptyset\}$ yields the desired construction. Fix $n\in G$. Since $\lambda^{\prime}B^{\ast}$ is a Polish space and $\widetilde{F}_{\lambda^{\prime}}$ is $\mathscr{A}_{R_{\lambda^{\prime}}}$-measurable (hence also Effros $\mathscr{A}_{R_{\lambda^{\prime}}}$-measurable, see Lemma 18.2 of [1]), the Kuratowski-Ryll-Nardzewski selection theorem (see pg. 600, [1]) implies the existence of an $\mathscr{A}$-measurable $h_{n}:D_{n}\longrightarrow E^{\ast}$ with $h_{n}(a)\in F(a)$ for each $a\in D_{n}$. Allowing $n$ to vary, we may define $h$ by setting $h|_{D_{n}}=h_{n}$ for each $n\in G$. ∎
Lemma 2.
Let $F$ be $\mathscr{A}$-measurable. Then, for every $x\in E$, the function
$a\longmapsto\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle$
is $\mathscr{A}$-measurable.
Proof.
Let $\{f_{n}:n\in\mathbb{N}\}$ be as in Lemma 1. Since the map $x^{\ast}\longmapsto\langle{x,x^{\ast}}\rangle$ is $\sigma(E^{\ast},E)$-continuous for each $x\in E$, it follows that, for each $x\in E$,
$\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle=\sup_{n\in\mathbb{N}}\langle{x,f_{n}(a)}\rangle,$
representing $a\longmapsto\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle$ as a countable supremum of $\mathscr{A}$-measurable functions. ∎
A.3. Statement of the Result
Theorem 6.
Let $F$ be an $\mathscr{A}$-measurable correspondence with non-empty $\sigma(E^{\ast},E)$-closed values. Suppose $S^{1}(F)\neq\emptyset$. For all $x\in E$, we have that
$\sup_{x^{\ast}\in\int_{A}F(a)\mu(da)}\langle{x,x^{\ast}}\rangle=\int_{A}\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle\mu(da).$
Note that Lemma 2 implies that $a\longmapsto\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle$ is $\mathscr{A}$-measurable, making the integral in Theorem 6 above well-defined.
Let us state a corollary to Theorem 6, which is more directly applicable to our situation with risk measures than Theorem 6. Rather than being interested in the supremum of $\langle{x,x^{\ast}}\rangle$ over $x^{\ast}\in\int_{A}F(a)\mu(da)$, we are interested in the supremum of a larger set $C\supseteq\int_{A}F(a)\mu(da)$.
Theorem 7.
Let $F$ be an $\mathscr{A}$-measurable correspondence with non-empty $\sigma(E^{\ast},E)$-closed values. Suppose $S^{1}(F)\neq\emptyset$, and let $C\subseteq E^{\ast}$ be such that $C=\overline{\int_{A}F(a)\mu(da)}^{\mathscr{T}}$ for some topology $\mathscr{T}$ finer than $\sigma(E^{\ast},E)$. For all $x\in E$, we have that
$\sup_{x^{\ast}\in C}\langle{x,x^{\ast}}\rangle=\int_{A}\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle\mu(da).$
Proof.
The claim is a trivial joint consequence of Theorem 6 and $\mathscr{T}$-continuity of the map $x^{\ast}\longmapsto\langle{x,x^{\ast}}\rangle$. ∎
A.4. Proof of Theorem 6
Proof of Theorem 6.
Let $\Xi(a)=\sup_{x^{\ast}\in F(a)}\langle{x,x^{\ast}}\rangle$. If the claim were false, there would exist $\beta<\int_{A}\Xi(a)\mu(da)$ such that $\beta>\left\langle{x,\int_{A}g(a)\mu(da)}\right\rangle$ for each $g\in S^{1}(F)$. Fix an arbitrary $f_{0}\in S^{1}(F)$ for later.
Let $(C_{n})_{n=1}^{\infty}\subseteq\mathscr{A}$ so $A\setminus C_{n}=\left\{\Xi\leq\widetilde{c}_{n}\right\}$ for some decreasing sequence $(\widetilde{c}_{n})_{n=1}^{\infty}\subseteq(-\infty,0)$ with $(\widetilde{c}_{n})_{n=1}^{\infty}\downarrow$-$\infty$ and
$\left|{\int_{\left\{\Xi\leq\widetilde{c}_{n}\right\}}}\langle{x,f_{0}(a)}\rangle\mu(da)\right|\leq\frac{1}{n}$
for all $n$. We may find a sequence of simple functions $(\Xi_{n})_{n=1}^{\infty}$ so $C_{n}\subseteq\{\Xi_{n}\leq\Xi\}$, $A\setminus C_{n}\subseteq\{\Xi_{n}\leq 0\}$, and there exists a small $\delta>0$ so $\int_{A}\Xi_{n}(a)\mu(da)>\beta+\delta$ for sufficiently large $n$ (say, $n\geq n_{1}$).
For any $n\in\mathbb{N}$, define a correspondence $G_{n}:A\longrightarrow 2^{E^{\ast}}$ by setting
$G_{n}(a)=\left\{x^{\ast}:\langle{x,x^{\ast}}\rangle\geq\Xi_{n}(a)-\frac{1}{n}\right\}\cap F(a)$
for $a\in C_{n}$, and $G_{n}(a)=\{f_{0}(a)\}$ for $a\in A\setminus C_{n}$. It is easy to see that $G_{n}$ is $\mathscr{A}$-measurable: if $U\subseteq E^{\ast}$ is $\sigma(E^{\ast},E)$-closed, then
$G^{-1}_{n}(U)=\left(\left(\bigcup_{\alpha\in\Xi_{n}(A)}F^{-1}\left(U\cap\left\{x^{\ast}:\langle{x,x^{\ast}}\rangle\geq\alpha-\frac{1}{n}\right\}\right)\cap\left\{\Xi_{n}=\alpha\right\}\right)\cap C_{n}\right)$
$\cup\left(f_{0}^{-1}(U)\cap(A\setminus C_{n})\right)$
which is evidently $\mathscr{A}$-measurable.
Fix an arbitrary $n\in\mathbb{N}$. By the same argument as in the proof of Lemma 1, there exists an $\mathscr{A}$-measurable $h_{n}:A\longrightarrow E^{\ast}$ such that $h_{n}(a)\in G_{n}(a)$ for each $a\in A$. For $m\in\mathbb{N}$, define $B^{n,m}=h_{n}^{-1}(mB^{\ast})\in\mathscr{A}$, and set $h_{n,m}=\mathbf{1}_{B^{n,m}}h_{n}+\mathbf{1}_{A\setminus B^{n,m}}f_{0}$. Clearly, $h_{n,m}\in S^{1}(F)$. Notice that
$\left\langle{x,\int_{A}h_{n,m}(a)\mu(da)}\right\rangle=\left\langle{x,\int_{B^{n,m}}h_{n}(a)\mu(da)}\right\rangle+\left\langle{x,\int_{A\setminus B^{n,m}}f_{0}(a)\mu(da)}\right\rangle$
$\left\langle{x,\int_{B^{n,m}\cap C_{n}}}h_{n}(a)\mu(da)\right\rangle+\left\langle{x,\int_{B^{n,m}\cap(A\setminus C_{n})}}h_{n}(a)\mu(da)\right\rangle+\left\langle{x,\int_{A\setminus B^{n,m}}f_{0}(a)\mu(da)}\right\rangle$
$\geq\int_{B^{n,m}\cap C_{n}}\left(\Xi_{n}(a)-\frac{1}{n}\right)\mu(da)+\left\langle{x,\int_{\left(B^{n,m}\cap(A\setminus C_{n})\right)\cup\left(A\setminus B^{n,m}\right)}f_{0}(a)\mu(da)}\right\rangle.$
Taking $m\to\infty$ in the last expression, we get $\int_{C_{n}}\left(\Xi_{n}(a)-\frac{1}{n}\right)\mu(da)+\int_{A\setminus C_{n}}\langle{x,f_{0}(a)}\rangle\mu(da)$, and from this limiting expression we obtain the chain of inequalities
$\int_{C_{n}}\left(\Xi_{n}(a)-\frac{1}{n}\right)\mu(da)+\int_{A\setminus C_{n}}\langle{x,f_{0}(a)}\rangle\mu(da)\geq\int_{C_{n}}\left(\Xi_{n}(a)-\frac{1}{n}\right)\mu(da)-\frac{\mu(A\setminus C_{n})}{n}$
$\int_{C_{n}}\Xi_{n}(a)\mu(da)-\frac{\mu(A)}{n}\geq\int_{A}\Xi_{n}(a)\mu(da)-\frac{\mu(A)}{n}.$
So, for any $\varepsilon>0$, there exists $f_{n}\in S^{1}(F)$ (more precisely, we can take $f_{n}=h_{n,m}$ for sufficiently large $m$) so that
$\left\langle{x,\int_{A}f_{n}(a)\mu(da)}\right\rangle+\varepsilon\geq\int_{A}\Xi_{n}(a)\mu(da)-\frac{\mu(A)}{n}.$
Taking $\varepsilon=\frac{1}{n}$, it follows that there exists $f_{n}\in S^{1}(F)$ so that
$\left\langle{x,\int_{A}f_{n}(a)\mu(da)}\right\rangle\geq\int_{A}\Xi_{n}(a)\mu(da)-\frac{\mu(A)+1}{n}.$ (9)
Using the arbitrariness of $n\in\mathbb{N}$, we have a sequence $(f_{n})_{n=1}^{\infty}\subseteq S^{1}(F)$ so that each $f_{n}$ satisfies (9).
Let $n\geq n_{1}$ be sufficiently large so that $\delta-\frac{\mu(A)+1}{n}>0$. Then
$\left\langle{x,\int_{A}f_{n}(a)\mu(da)}\right\rangle\geq\int_{A}\Xi_{n}(a)\mu(da)-\frac{\mu(A)+1}{n}>\beta+\delta-\frac{\mu(A)+1}{n}>\beta.$
Hence, there exists $f\in S^{1}(F)$ (namely, $f=f_{n}$ for some $n\geq n_{1}$, as above) so that $\left\langle{x,\int_{A}f_{n}(a)\mu(da)}\right\rangle>\beta$, a contradiction, as $\beta$ is assumed to satisfy the reverse inequality. ∎
Appendix B Proof of Theorem 1
In this section, we prove Theorem 1. Various tools are employed, including results from Appendix A, and some results about acceptance sets (see §B.1 below).
B.1. Acceptance Sets
Given a risk measure $\varrho$, the acceptance set $\mathfrak{A}(\varrho)$ is defined by
$\mathfrak{A}(\varrho)=\{\mathcal{X}:\varrho(\mathcal{X})\leq 0\}.$
The Fatou property implies that $\mathfrak{A}(\varrho)$ is $\sigma(L^{\infty},L^{1})$-closed.
B.1.1. Characterizing the Acceptance Set
In this subsection, we provide a characterization of the acceptance set of the integral infimal convolution of $(\varrho_{a})_{a\in A}$, in terms of the closure of a certain Aumann integral.
Recall the Aumann integral from Definition 10 in Appendix A, which we reproduce here in a slightly less abstract setting. Given a correspondence $F:A\longrightarrow 2^{L^{\infty}(\mathbb{P})}$, an integrable selector of $F$ is an $\mathscr{A}$-measurable Gelfand integrable function $(X_{a})_{a\in A}\in\left(L^{\infty}(\mathbb{P})\right)^{A}$ such that $X_{a}\in F(a)$ for $\mu$-a.e. $a\in A$. The set of all integrable selectors of $F$ is denoted $S^{1}(F)$. The Aumann integral $\int_{A}F(a)\mu(da)$ of $F$ is defined as
$\int_{A}F(a)\mu(da)=\left\{\int_{A}X_{a}\mu(da):(X_{a})_{a\in A}\in S^{1}(F)\right\}.$
Theorem 8.
Suppose $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is globally finite. Then, the acceptance set $\mathfrak{A}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)$ of $\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)$ is the $L^{\infty}(\mathbb{P})$-closure of the Aumann integral $\int_{A}\mathfrak{A}(\varrho_{a})\mu(da)$.
Proof.
We follow the same idea as the proof of (Theorem 4.1, [20]), replacing finite Minkowski sums with Aumann integrals.
Clearly, $\int_{A}\mathfrak{A}(\varrho_{a})\mu(da)\subseteq\mathfrak{A}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)$. Since the integral infimal convolution is monotone and cash additive, we may apply the same argument as (Lemma 4.3, [11]) to obtain that $\mathfrak{A}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)$ is $L^{\infty}(\mathbb{P})$-closed, implying $\overline{\int_{A}\mathfrak{A}(\varrho_{a})\mu(da)}^{L^{\infty}}\subseteq\mathfrak{A}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)$. Thus, it suffices to show the reverse inclusion.
Let $\mathcal{X}\in\mathfrak{A}\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)$; denote $w=\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})\leq 0$. By cash additivity, $\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)\left(\mathcal{X}-w\right)=0$. Thus, there exists a sequence $\left((X^{n}_{a})_{a\in A}\right)_{n=1}^{\infty}\subseteq\mathbb{A}(\mathcal{X}-w)$ such that
$\lim_{n\to\infty}\int_{A}\varrho_{a}\left(X^{n}_{a}\right)\mu(da)=0.$
Let $Y^{n}_{a}=X^{n}_{a}-\varrho_{a}(X^{n}_{a})$. Clearly, $(Y^{n}_{a})_{a\in A}\in\mathfrak{A}(\varrho_{a})$, implying that $\int_{A}Y^{n}_{a}\mu(da)\in\int_{A}\mathfrak{A}(\varrho_{a})\mu(da)$. Thus, since the $L^{\infty}(\mathbb{P})$-limit of $\left(\int_{A}Y^{n}_{a}\mu(da)\right)_{n=1}^{\infty}$ is $\mathcal{X}-w$, it follows that $\mathcal{X}-w\in\overline{\int_{A}\mathfrak{A}(\varrho_{a})\mu(da)}^{L^{\infty}}$. Since $w\leq 0$, it follows that $\mathcal{X}\in\overline{\int_{A}\mathfrak{A}(\varrho_{a})\mu(da)}^{L^{\infty}}$, as desired. ∎
B.1.2. Representing the Dual via Acceptance Sets
In this subsection, we state a known result connecting acceptance sets to convex conjugates. This allows us to apply our results on the correspondence $a\longmapsto\mathfrak{A}(\varrho_{a})$ to dual representations.
Lemma 3.
Let $\varrho$ be any risk measure. Then, for any $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$,
$\varrho^{\ast}(\mathbb{Q})=\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X}).$
Proof.
Fix $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$; clearly, $\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})\leq\varrho^{\ast}(\mathbb{Q})$. By cash additivity,
$\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})\leq\varrho^{\ast}(\mathbb{Q})=\sup_{\mathcal{X}\in L^{\infty}(\mathbb{P})}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X})-\varrho(\mathcal{X})\right)$
$\sup_{\mathcal{X}\in L^{\infty}(\mathbb{P})}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X}-\varrho(X))-\varrho(\mathcal{X}-\varrho(X))\right)=\sup_{\mathcal{X}\in\{\varrho=0\}}\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)\leq\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$
implying that $\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})\leq\varrho^{\ast}(\mathbb{Q})\leq\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$, showing that $\varrho^{\ast}(\mathbb{Q})=\sup_{\mathcal{X}\in\mathfrak{A}(\varrho)}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$. Since $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$ was arbitrary, this proves the claim. ∎
B.1.3. Measurability of the Acceptance Set Correspondence
In this subsection, we establish that the acceptance set correspondence $a\longmapsto\mathfrak{A}(\varrho_{a})$ is $\mathscr{A}$-measurable in the sense of Definition 11 from Appendix A. We assume in this subsection that $(\varrho_{a})_{a\in A}$ denotes an $\mathscr{A}$-measurable indexed collection of risk measures such that each $\varrho_{a}$ has the Lebesgue property.
Lemma 4.
Let $U\subseteq L^{\infty}(\mathbb{P})$ be $\sigma(L^{\infty},L^{1})$-closed. Then
$\left\{a\in A:U\cap\mathfrak{A}(\varrho_{a})\neq\emptyset\right\}\in\mathscr{A}.$
Proof.
It is no loss of generality to assume that $U$ is bounded in $L^{\infty}(\mathbb{P})$, since one can write $U$ as a countable union of closed and $L^{\infty}(\mathbb{P})$-bounded sets. Furthermore, we may assume that $U\neq\emptyset$ (if $U=\emptyset$, the claim would be trivial).
We claim that
$\left\{a\in A:U\cap\mathfrak{A}(\varrho_{a})\neq\emptyset\right\}=\left\{a\in A:\inf_{\mathcal{Y}\in U}\varrho_{a}(\mathcal{Y})\leq 0\right\}.$ (10)
Clearly, $\left\{a\in A:U\cap\mathfrak{A}(\varrho_{a})\neq\emptyset\right\}\subseteq\left\{a\in A:\inf_{\mathcal{Y}\in U}\varrho_{a}(\mathcal{Y})\leq 0\right\}$. Thus, it suffices to show the reverse inclusion. If $\inf_{\mathcal{Y}\in U}\varrho_{a}(\mathcal{Y})\leq 0$, there exists $(\mathcal{Y}^{n})_{n=1}^{\infty}\subseteq U$ such that
$\varrho_{a}\left(\mathcal{Y}^{n}\right)\leq\frac{1}{n}.$
Using the Banach-Alaoglu theorem, $L^{\infty}(\mathbb{P})$-boundedness of $U$, and $\sigma(L^{\infty},L^{1})$-closedness of $U$, we may find a subsequence $(n_{k})_{k=1}^{\infty}$ such that $(\mathcal{Y}^{n_{k}})_{k=1}^{\infty}$ converges to some $\mathcal{Z}\in U$ in $\sigma(L^{\infty},L^{1})$. For each $n$, there exists $k_{0}$ such that $k\geq k_{0}$ implies $\mathcal{Y}^{n_{k}}\in\{\varrho_{a}\leq\frac{1}{n}\}$. The Fatou property implies the set $\left\{\varrho_{a}\leq\frac{1}{n}\right\}$ is $\sigma(L^{\infty},L^{1})$-closed, and we therefore have that $\mathcal{Z}\in\left\{\varrho_{a}\leq\frac{1}{n}\right\}$ for each $n$. Thus,
$\mathcal{Z}\in\bigcap_{n\in\mathbb{N}}\left\{\varrho_{a}\leq\frac{1}{n}\right\}=\{\varrho_{a}\leq 0\}.$
By the above argument, there exists $\mathcal{Z}\in U$ with $\varrho_{a}(\mathcal{Z})\leq 0$, implying that $a\in\left\{b\in A:U\cap\mathfrak{A}(\varrho_{b})\neq\emptyset\right\}$, as desired.
As a consequence of (10), it suffices to show that $a\longmapsto\inf_{\mathcal{Y}\in U}\varrho_{a}(\mathcal{Y})$ is $\mathscr{A}$-measurable. Let $V\subseteq U$ be a countable dense set for the topology $\tau_{L^{0}}$ of convergence in probability restricted to $U$ (such a set exists, since $(\Omega,\mathscr{F},\mathbb{P})$ is separable). We claim that
$\inf_{\mathcal{Y}\in U}\varrho_{a}(\mathcal{Y})=\inf_{\mathcal{Y}\in V}\varrho_{a}(\mathcal{Y}),$ (11)
for all $a\in A$, which would prove the claim, since $a\longmapsto\inf_{\mathcal{Y}\in U}\varrho_{a}(\mathcal{Y})$ would be a countable infimum of $\mathscr{A}$-measurable functions. Since $\varrho_{a}$ has the Lebesgue property, $U$ is $L^{\infty}(\mathbb{P})$-bounded, and $\overline{V}^{\tau_{L^{0}}}=U$, (11) holds. ∎
B.2. The Proof
Proof of Theorem 1.
Recall that
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})=\sup_{\mathcal{X}\in L^{\infty}(\mathbb{P})}\left(\mathbb{E}^{\mathbb{Q}}(\mathcal{X})-\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)(\mathcal{X})\right)$
for each $\mathbb{Q}\in\mathscr{M}_{\mathbb{P}}$. By Lemma 3 and Theorem 8,
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})=\sup_{\mathcal{X}\in\overline{\int_{A}\mathfrak{A}(\varrho_{a})}^{L^{\infty}}}\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right).$ (12)
We claim that $S^{1}(F)\neq\emptyset$ ($S^{1}(F)$ is defined in §B.1.1), where $F$ is the correspondence $F=\left(a\longmapsto\mathfrak{A}(\varrho_{a})\right)$. Define $(X_{a})_{a\in A}\in\left(L^{\infty}(\mathbb{P})\right)^{A}$ by setting
$X_{a}=-\varrho_{a}(0).$
It is easy to see that $(X_{a})_{a\in A}$ is $\mathscr{A}$-measurable; furthermore, since $\int_{A}|{\varrho_{a}(0)}|\mu(da)<\infty$, $(X_{a})_{a\in A}$ is Gelfand integrable. For each $a\in A$, we have that
$\varrho_{a}(X_{a})=\varrho_{a}(-\varrho_{a}(0))=\varrho_{a}(0)+(-\varrho_{a}(0))=0\leq 0$
by cash additivity. Thus, $X_{a}\in F(a)$ for each $a\in A$. These facts together imply that $(X_{a})_{a\in A}\in S^{1}(F)$, showing that $S^{1}(F)\neq\emptyset$.
By Lemma 4, $F$ is $\mathscr{A}$-measurable in the sense of Definition 11 from Appendix A. Thus, since $S^{1}(F)\neq\emptyset$, the preconditions for Theorem 6 and Theorem 7 are met. Noting that the norm topology on $L^{\infty}(\mathbb{P})$ is finer than $\sigma(L^{\infty},L^{1})$, Theorem 7 and (12) imply that
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})=\int_{A}\sup_{\mathcal{X}\in F(a)}\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)\mu(da),$
where we note that $a\longmapsto\sup_{\mathcal{X}\in F(a)}\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)$ is $\mathscr{A}$-measurable, by virtue of Lemma 2. By Lemma 3, we have that $\varrho^{\ast}_{a}(\mathbb{Q})=\sup_{\mathcal{X}\in F(a)}\mathbb{E}^{\mathbb{Q}}\left(\mathcal{X}\right)$ for each $a\in A$, which implies that $a\longmapsto\varrho^{\ast}(\mathbb{Q})$ is $\mathscr{A}$-measurable, and that
$\left(\operatorname{\Box}_{a\in A}\varrho_{a}\mu(da)\right)^{\ast}(\mathbb{Q})=\int_{A}\varrho^{\ast}_{a}(\mathbb{Q})\mu(da),$
as desired. ∎
Appendix C Proof of Theorem 4
In this section, we prove the second part of Theorem 4, showing that under certain circumstances the risk sharing problem is not well-posed.
C.1. Lemmata
Before we proceed, remark that if $\varrho$ is a coherent risk measure with the Lebesgue property, the dual representation of $\varrho$ is of the form
$\varrho(\mathcal{X})=\sup_{\mathbb{Q}\in\{\varrho^{\ast}<\infty\}}\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$
for any $\mathcal{X}\in L^{\infty}(\mathbb{P})$, where the supremum is attained for any fixed $\mathcal{X}$ because of the dual characterization of the Lebesgue property due to Jouini, Schachermayer, and Touzi (see Theorem 2.4, [21]). In a previous context, we did not assume the Lebesgue property, and hence we could only select a measure $\mathbb{Q}\in\{\varrho^{\ast}<\infty\}$ under which the expectation $\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$ approximates, rather than equals, $\varrho(\mathcal{X})$ (see the proof of Proposition 3). In the situations below, all risk measures have the Lebesgue property, and so for any fixed $\mathcal{X}$ we may always choose a probability measure $\mathbb{Q}\in\{\varrho^{\ast}<\infty\}$ so that $\varrho(\mathcal{X})=\mathbb{E}^{\mathbb{Q}}(\mathcal{X})$.
The following lemma is related to the behavior of expected shortfall as a function of quantile level.
Lemma 5.
Fix a random variable $\mathcal{X}\in L^{\infty}(\mathbb{P})$. If $\alpha,\alpha^{\prime},\alpha^{\prime\prime}\in(0,1)$, $\alpha^{\prime\prime}<\alpha^{\prime}<\alpha$, and $\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})=\mathrm{ES}^{\alpha^{\prime}}_{\mathbb{P}}(\mathcal{X})$, then $\mathrm{ES}^{\alpha^{\prime\prime}}_{\mathbb{P}}(\mathcal{X})=\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})$.
Proof.
Suppose, for the sake of contradiction, that $\mathrm{ES}^{\alpha^{\prime\prime}}_{\mathbb{P}}(\mathcal{X})>\delta+\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})$ for some $\delta>0$. Write $\frac{1}{\alpha^{\prime}}$ as a convex combination $\frac{1}{\alpha^{\prime}}=\lambda\frac{1}{\alpha}+(1-\lambda)\frac{1}{\alpha^{\prime\prime}}$ for some $\lambda\in(0,1)$. We may find probability measures $\mathbb{Q}_{1},\mathbb{Q}_{2}\ll\mathbb{P}$ with $\frac{d\mathbb{Q}_{1}}{d\mathbb{P}}\leq\frac{1}{\alpha}$, $\frac{d\mathbb{Q}_{2}}{d\mathbb{P}}\leq\frac{1}{\alpha^{\prime\prime}}$, $\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})=\mathbb{E}^{\mathbb{Q}_{1}}(\mathcal{X})$, and $\mathrm{ES}^{\alpha^{\prime\prime}}_{\mathbb{P}}(\mathcal{X})=\mathbb{E}^{\mathbb{Q}_{2}}(\mathcal{X})$. Setting $\mathbb{Q}_{3}=\lambda\mathbb{Q}_{1}+(1-\lambda)\mathbb{Q}_{2}$, we have that $\frac{d\mathbb{Q}_{3}}{d\mathbb{P}}\leq\lambda\frac{1}{\alpha}+(1-\lambda)\frac{1}{\alpha^{\prime\prime}}=\frac{1}{\alpha^{\prime}}$, and hence $\mathbb{E}^{\mathbb{Q}_{3}}(\mathcal{X})\leq\mathrm{ES}^{\alpha^{\prime}}_{\mathbb{P}}(\mathcal{X})$. But $\mathbb{E}^{\mathbb{Q}_{3}}(\mathcal{X})=\lambda\mathbb{E}^{\mathbb{Q}_{1}}(\mathcal{X})+(1-\lambda)\mathbb{E}^{\mathbb{Q}_{2}}(\mathcal{X})=\lambda\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})+(1-\lambda)\mathrm{ES}^{\alpha^{\prime\prime}}_{\mathbb{P}}(\mathcal{X})>(1-\lambda)\delta+\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})>\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})=\mathrm{ES}^{\alpha^{\prime}}_{\mathbb{P}}(\mathcal{X})$, a contradiction. ∎
As a corollary to the previous lemma, we have the following.
Lemma 6.
Fix a random variable $\mathcal{X}\in L^{\infty}(\mathbb{P})$. If $\alpha,\alpha^{\prime}\in(0,1)$, $\alpha^{\prime}<\alpha$, and $\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})=\mathrm{ES}^{\alpha^{\prime}}_{\mathbb{P}}(\mathcal{X})$, then $\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})=\mathbb{P}-\mathrm{essup}\mathcal{X}$.
Proof.
As $\beta\to 0$, $\mathrm{ES}^{\beta}_{\mathbb{P}}(\mathcal{X})\longrightarrow\mathbb{P}-\mathrm{essup}\mathcal{X}$, and applying the previous lemma yields the claim, as $\mathrm{ES}^{\beta}_{\mathbb{P}}(\mathcal{X})=\mathrm{ES}^{\alpha}_{\mathbb{P}}(\mathcal{X})$ for sufficiently small $\beta$. ∎
The previous results of this subsection are entirely about expected shortfall, but they generalize to inflations of a large class of risk measures.
Lemma 7.
Let $\varrho$ be a risk measure with the Lebesgue property and the representation (7), and fix a random variable $\mathcal{X}\in L^{\infty}(\mathbb{P})$. If $\gamma,\gamma^{\prime},\gamma^{\prime\prime}\in[1,\infty)$, $\gamma^{\prime\prime}\geq\gamma^{\prime}>\gamma$, and $\widetilde{\varrho}_{\gamma}(\mathcal{X})=\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})$, then $\widetilde{\varrho}_{\gamma^{\prime\prime}}(\mathcal{X})=\widetilde{\varrho}_{\gamma}(\mathcal{X})$.
Proof.
Define $q(x)=\frac{1}{x}$. Fix some $\widetilde{\gamma}\in(\gamma,\gamma^{\prime})$, and write $\widetilde{\gamma}=\frac{\gamma}{1-\delta}$ for $0<\delta<1$. Find a probability measure $\mathbb{Q}\in\mathscr{Q}(\varrho)$ so there exists a probability $\mathbb{Q}_{1}\ll\mathbb{Q}$ with $\frac{d\mathbb{Q}_{1}}{d\mathbb{Q}}\leq\gamma$ and $\mathbb{E}^{\mathbb{Q}_{1}}(\mathcal{X})=\widetilde{\varrho}_{\gamma}(\mathcal{X})$. In particular, $\widetilde{\varrho}_{\gamma}(\mathcal{X})=\mathrm{ES}^{q(\gamma)}_{\mathbb{Q}}(\mathcal{X})$.
Set $\mathbb{Q}_{2}=\delta\mathbb{P}+(1-\delta)\mathbb{Q}\sim\mathbb{P}$. Define $\mathcal{Z}$ as $\frac{d\mathbb{Q}_{1}}{d\mathbb{Q}}$ on $\left\{\frac{d\mathbb{Q}}{d\mathbb{P}}>0\right\}$, and zero elsewhere. For any $B\in\mathscr{F}$,
$\mathbb{Q}_{1}(B)\leq\int_{B}\frac{1}{1-\delta}\mathcal{Z}d\mathbb{Q}_{2}$
and hence
$\frac{d\mathbb{Q}_{1}}{d\mathbb{Q}_{2}}\leq\frac{1}{1-\delta}\mathcal{Z}\leq\frac{\gamma}{1-\delta}=\widetilde{\gamma}$.
It follows that $\mathrm{ES}^{q(\widetilde{\gamma})}_{\mathbb{Q}_{2}}(\mathcal{X})=\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})$, and hence $\mathrm{ES}^{q(\widetilde{\gamma})}_{\mathbb{Q}_{2}}(\mathcal{X})=\mathrm{ES}^{q(\gamma^{\prime})}_{\mathbb{Q}_{2}}(\mathcal{X})$. By Lemma 6, we have that $\mathrm{ES}^{q(\gamma^{\prime})}_{\mathbb{Q}_{2}}(\mathcal{X})=\mathbb{Q}_{2}-\mathrm{essup}\mathcal{X}=\mathbb{P}-\mathrm{essup}\mathcal{X}$, so $\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})=\mathbb{P}-\mathrm{essup}\mathcal{X}$. As $\widetilde{\varrho}_{\gamma}(\mathcal{X})=\widetilde{\varrho}_{\gamma^{\prime}}(\mathcal{X})$ and
$\widetilde{\varrho}_{\gamma}(\mathcal{X})\leq\widetilde{\varrho}_{\gamma^{\prime\prime}}(\mathcal{X})\leq\mathbb{P}-\mathrm{essup}\mathcal{X}$
this finishes the proof. ∎
C.2. The Proof
Proof of Theorem 4.
We prove the second assertion. Assume, for contradiction, that $(X_{a})_{a\in A}$ is optimal and the preconditions of the second assertion are satisfied.
For each $n\in\mathbb{N}$, define
$B_{n}=\left\{a\in A:\gamma_{a}\geq\Gamma+\frac{1}{n}\right\},$
$C_{n}=\left\{a\in A:\gamma_{a}<\Gamma+\frac{1}{n}\right\}.$
As $(X_{a})_{a\in A}$ is optimal, $\widetilde{\varrho}_{\Gamma}(\mathcal{X})=\int_{A}\widetilde{\varrho}_{\gamma_{a}}(X_{a})\mu(da)$ by Theorem 3. Hence
$\widetilde{\varrho}_{\Gamma}(\mathcal{X})=\int_{B_{n}}\widetilde{\varrho}_{\gamma_{a}}(X_{a})\mu(da)+\int_{C_{n}}\widetilde{\varrho}_{\gamma_{a}}(X_{a})\mu(da)$
$\geq\widetilde{\varrho}_{\Gamma}\left(\int_{C_{n}}X_{a}\mu(da)\right)+\widetilde{\varrho}_{\Gamma+\frac{1}{n}}\left(\int_{B_{n}}X_{a}\mu(da)\right).$
Suppose, for some $n\in\mathbb{N}$, that $\widetilde{\varrho}_{\Gamma+\frac{1}{n}}\left(\int_{B_{n}}X_{a}\mu(da)\right)>\widetilde{\varrho}_{\Gamma}\left(\int_{B_{n}}X_{a}\mu(da)\right)$. Then
$\widetilde{\varrho}_{\Gamma}(\mathcal{X})\geq\widetilde{\varrho}_{\Gamma}\left(\int_{C_{n}}X_{a}\mu(da)\right)+\widetilde{\varrho}_{\Gamma+\frac{1}{n}}\left(\int_{B_{n}}X_{a}\mu(da)\right)$
$\widetilde{\varrho}_{\Gamma}\left(\int_{C_{n}}X_{a}\mu(da)\right)+\widetilde{\varrho}_{\Gamma}\left(\int_{B_{n}}X_{a}\mu(da)\right)\geq\widetilde{\varrho}_{\Gamma}(\mathcal{X})$
by the subadditivity of coherent risk measures, a contradiction. So, for all $n\in\mathbb{N}$, $\widetilde{\varrho}_{\Gamma+\frac{1}{n}}\left(\int_{B_{n}}X_{a}\mu(da)\right)=\widetilde{\varrho}_{\Gamma}\left(\int_{B_{n}}X_{a}\mu(da)\right)$. Lemma 7 implies
$\widetilde{\varrho}_{\Gamma}\left(\int_{B_{n}}X_{a}\mu(da)\right)=\widetilde{\varrho}_{\Gamma^{\prime}}\left(\int_{B_{n}}X_{a}\mu(da)\right)$ (13)
for all sufficiently large $n\in\mathbb{N}$ (large enough so that $\Gamma^{\prime}\geq\Gamma+\frac{1}{n}$).
Define a linear operator $G:L^{1}(\mathbb{P})\longrightarrow L^{1}(\mu)$ by $G(\mathcal{Y})=\mathbb{E}^{\mathbb{P}}(X_{a}\mathcal{Y})$. It is easy to see that $G$ is weak to weak continuous. Hence, if $D$ denotes the unit ball of $L^{\infty}(\mathbb{P})$, the Dunford-Pettis theorem implies $G(D)$ is uniformly $\mu$-integrable. So, for any $\varepsilon>0$, there exists $\delta>0$ so $\mu(E)<\delta$ implies
$\left\|{\int_{E}X_{a}\mu(da)}\right\|_{L^{1}(\mathbb{P})}=\sup_{\mathcal{Y}\in D}\left|\int_{E}\mathbb{E}^{\mathbb{P}}(X_{a}\mathcal{Y})\mu(da)\right|<\varepsilon.$
As $n\to\infty$, $\mu(C_{n})\to 0$, and in particular $\mu(C_{n})<\delta$ for large enough $n$, so that
$\left\|{\mathcal{X}-\int_{B_{n}}X_{a}\mu(da)}\right\|_{L^{1}(\mathbb{P})}=\left\|{\int_{C_{n}}X_{a}\mu(da)}\right\|_{L^{1}(\mathbb{P})}<\varepsilon$
for large enough $n$. As $\varepsilon>0$ was arbitrary, we have that $\left(\int_{B_{n}}X_{a}\mu(da)\right)_{n=1}^{\infty}$ converges to $\mathcal{X}$ in $L^{1}(\mathbb{P})$, and hence also in probability. Further, $\left(\int_{B_{n}}X_{a}\mu(da)\right)_{n=1}^{\infty}$ evidently converges in $\sigma(L^{\infty},L^{1})$ to $\mathcal{X}$, so the Banach-Steinhaus theorem implies $\left(\int_{B_{n}}X_{a}\mu(da)\right)_{n=1}^{\infty}$ is bounded in $L^{\infty}$. We can thus apply the Lebesgue property of $\widetilde{\varrho}_{\Gamma}$ and $\widetilde{\varrho}_{\Gamma^{\prime}}$ to obtain
$\widetilde{\varrho}_{\Gamma}\left(\mathcal{X}\right)=\widetilde{\varrho}_{\Gamma^{\prime}}\left(\mathcal{X}\right)$
by taking $n\to\infty$ in (13), contradicting the assumption $\widetilde{\varrho}_{\Gamma}\left(\mathcal{X}\right)\neq\widetilde{\varrho}_{\Gamma^{\prime}}\left(\mathcal{X}\right)$. ∎